### v1

# The GenAI Generation: Student Views of Awareness, Preparedness, and Concern

Abstract
Generative AI (GenAI) is revolutionizing education and workforce development, profoundly shaping how students learn, engage, and prepare for their future. Outpacing the development of uniform policies and structures, GenAI has heralded a unique era and given rise to the GenAI Generation: a cohort of students whose education has been increasingly shaped by the opportunities and challenges GenAI presents during its widespread adoption within society. This study examines our students’ perceptions of GenAI through a concise survey with optional open-ended questions, focusing on their awareness, preparedness, and concerns. Evaluation of more than 250 responses with more than 40% providing detailed qualitative feedback reveals a core dual sentiment: while most students express enthusiasm for GenAI, an even greater proportion voice a spectrum of concerns about ethics, job displacement, and the adequacy of educational structures given the highly transformative technology. These findings offer critical insights into how students view the potential and pitfalls of GenAI for future career impacts, with accompanying recommendations to guide educational institutions in navigating a future driven by GenAI.
1 Introduction
The rapid advancement of artificial intelligence (AI), particularly generative AI (GenAI), is reshaping the educational landscape. As a tool capable of autonomously producing human-like content, GenAI offers both opportunities and uncertainties for education. It can enhance creativity, streamline workflows, and support personalized learning, yet its broader impact remains unclear. As GenAI becomes more embedded in academic and professional contexts, understanding how students perceive its influence on future readiness is essential.
This study presents the development, dissemination, and evaluation of a survey examining student perspectives on GenAI through three core lenses: awareness, preparedness, and concern. By analyzing responses, this research provides actionable insights into how the GenAI Generation, students most affected by GenAI’s rapid integration, understand its role in shaping their academic and career paths.
Defining the GenAI Generation
The GenAI Revolution represents a paradigm shift in technology adoption, enhancing efficiency, creativity, and performance while introducing challenges such as ethical concerns and the need for responsible stewardship [14, 6, 19, 5]. As it relates to education, this revolution not only emphasizes a transformation in learning but also highlights the challenges students and teachers face during a phase of unclear uniform guidance [5, 6]. This study introduces the concept of the GenAI Generation to describe individuals most directly impacted by the Generative AI Revolution. The GenAI Generation is defined as the cohort of individuals actively undergoing developmental and educational growth/attainment during the rise of Generative AI, starting with the release of ChatGPT in November 2022. The GenAI Generation encompasses the experiences of students from high school through higher education, during the initial, transformative years of the GenAI Revolution, when the integration of AI tools is still experimental, unregulated, and highly influential. Unlike earlier generations who did not have these tools and later generations who may inherit more structured systems, this group faces unique challenges and opportunities surrounding uncertainty.
  • Developmental Context: Actively undergoing critical cognitive and social growth while navigating the GenAI Revolution.
  • Navigating Uncertainty: Early adopters, engaging with experimental, unregulated technologies without widely established policies.
  • High Adaptability: Balancing both the traditional and new competencies required to effectively interact and engage with GenAI in a rapidly changing environment.
  • Long-Term Impact Unknown: Using these tools during formative years without a clear understanding of potential long-term cognitive or social consequences.
Figure 1: Illustrates the relationships between key characteristics, emphasizing the ongoing feedback loop that drives development and adaptation amidst uncertainty.
Just as Google transformed how students accessed and retained information, leading to phenomena dubbed the “Google Effect” [31], and the rise of digital technologies sparked concerns about “digital dementia” [32], the GenAI Revolution represents another critical shift in the educational landscape. These past transformations demonstrate how new technologies can reshape cognitive processes, social behaviors, and academic growth, sometimes with unintended, long-term consequences [30, 7, 21]. The GenAI Generation, defined by its unique characteristics, is similarly at the forefront of this era’s challenges and opportunities. Evaluating this cohort is essential to understand how these tools influence learning and development during critical formative years, ensuring that we can build a foundation of guidance, policies, and research to support them and future generations effectively.
Awareness of GenAI’s Capabilities and Applications
For the purpose of this study, students’ awareness of GenAI means their familiarity with tools such as ChatGPT and their understanding of how these tools can be leveraged in educational and professional contexts. Research has revealed varying levels of awareness globally. One study found widespread awareness among students across 76 countries, with significant usage for academic tasks such as information retrieval and paraphrasing, though concerns about academic integrity and creativity persist [34]. Another study highlighted differences in engagement with GenAI based on demographic factors, underscoring the need for targeted educational support [13]. Research also shows that while GenAI is valued for idea generation and brainstorming, there are reservations regarding its use in completing assignments, emphasizing the need for ethical policies [4].
The transformative potential of GenAI in education has been widely discussed, particularly in its applications for personalized learning and adaptive tutoring systems, which support higher-order cognitive tasks such as problem-solving and critical analysis [9]. However, challenges like data privacy and instructor resistance remain significant barriers to effective integration. Other studies have explored how personalized learning and language support tools can aid students while also warning of potential over-reliance and its impact on critical thinking [20].
Current research on student awareness of GenAI’s impacts reveals perspectives shaped by demographic and geographic factors. These findings highlight the importance of adopting diverse approaches to educate students about GenAI while providing guidance on data privacy and ethical considerations.
Preparedness for an GenAI-Driven Workforce
For the purpose of this study, student preparedness refers to their confidence in navigating a dynamic, AI-driven job market and their perception of how well their education equips them for this future. Research highlights gaps in AI-specific training within higher education curricula, which leave students feeling under-prepared for the evolving workforce [1, 8]. Ethical and accuracy concerns are also noted as barriers to effective integration, further underscoring the importance of incorporating AI literacy and practical applications into academic programs to enhance student readiness [8]. Additionally, project-based and interdisciplinary learning approaches are recommended to prepare students to critically engage with AI tools and align their academic training with workforce demands [10].
Other studies propose professional development programs for educators to ensure that GenAI integration aligns with institutional goals and emphasizes critical digital literacy, enabling both students and faculty to navigate the complexities of GenAI in professional environments [25]. Personalized learning supported by GenAI is also identified as a potential way to improve academic and employability outcomes, though concerns about ethical use and critical thinking remain [15].
Students’ preparedness highlights the interdisciplinary importance of addressing GenAI in the classroom, ensuring clarity and intentionality in these discussions. Although this research does not examine faculty perspectives, it is critical to recognize that educators also require support and guidance to effectively integrate GenAI discussions and tools into their curricula.
Concerns About GenAI’s Potential Implications
For the purpose of this study, students’ concerns refer to their anxiety or worry about GenAI’s potential impact. Current research highlights ethical issues such as data privacy and bias, alongside practical challenges like overreliance and the risk of misinformation. Studies have noted strong reservations about using GenAI for completing assignments without disclosure as well as concerns about the uncritical acceptance of AI outputs [4, 2]. Additionally, research underscores how the misuse of GenAI could undermine academic integrity and originality, emphasizing the need for critical evaluation skills and clear institutional guidelines for responsible use [34, 2]. Balancing the benefits of automation with safeguards against biases and ethical risks remains a pressing concern.
Some researchers critique traditional assessment models, advocating for redesigned evaluations that prioritize originality and critical engagement [23]. Others warn that over-reliance on GenAI could diminish critical thinking and intellectual engagement, urging for its balanced integration into educational practices [20]. While GenAI tools have the potential to enhance advanced cognitive processes, concerns persist regarding their impact on foundational learning skills such as critical thinking and engagement [4, 34, 2, 23, 20].
This contrast highlights the urgent need for clear guidance on student use of GenAI across all educational levels. Strategies must be developed to ensure that these tools are employed ethically while fostering constructive critical engagement and safeguarding cognitive development.
Contributions of This Study
This study contributes to the ongoing dialogue about GenAI’s role in shaping students’ academic and professional paths. Centered on awareness, preparedness, and concern, the survey offers a structured view of the GenAI Generation, a group navigating rapid technological change during key years of education. Focusing on higher education, the analysis shows how students weigh GenAI’s potential to support creativity and critical thinking against concerns about overreliance, ethical risks, and reduced engagement. These insights emphasize the need for thoughtful strategies that support ethical and effective use of GenAI in education.
2 Methods
The survey targeted students within the GenAI Generation, focusing on GenAI’s perceived impact on their education and careers. Awareness, preparedness, and concern guided the development of eight main questions, three demographic items, and two open-ended prompts. Designed for completion in under five minutes [16], the survey began with an IRB-required disclosure and a screening question, followed by items progressing from general to sensitive topics [28]. Questions were refined through lab feedback, aligned with key concepts, and included internal validation to ensure consistency and quality.
A four-point, forced-choice Likert-type scale was used to assess attitudes, significance, intensity, and satisfaction. While omitting a neutral option is debated [17], this was an intentional design choice to encourage engagement and reduce bias [18]. The survey was administered across two cohorts, with full question text available in the supplemental materials.
The Test 2023 survey was distributed to Computer Science Master’s students via OMSCS (12,000 members) and MSCS (14,000 members) listservs. Invitations were sent on September 28 and 29, 2023, with responses collected through November 1. The Fall 2023 survey was conducted with support from the Office of Academic Effectiveness, which used stratified random sampling to select 5,655 undergraduates representative of the broader student population. Sampling quotas ensured balance by gender, race, citizenship, and class standing, with a 2% margin of error. The survey was open from November 2–23, 2023, with email reminders sent on November 2, 15, and 22.
Methods of Analysis
Survey responses were analyzed using correlation, clustering, and free-response evaluation to capture both quantitative and qualitative insights. These methods supported a systematic exploration of trends and perceptions across diverse student groups.
Data Processing: To ensure data integrity, only completed responses containing degree type information were included in the analysis. Degrees were categorized into three groups: Computer Science, Engineering, and Other, based on thematic alignment with disciplinary focus and methodologies. Multidisciplinary fields, such as Robotics and Computational Science and Engineering, were classified under Engineering and Computer Science, respectively, to maintain consistency. This classification enabled cross-disciplinary comparisons and analysis of AI’s impact on skills by degree category and education level.
For the Fall 2023 cohort, two non-undergraduate entries, likely from dual-enrolled students, were excluded to maintain a focus on undergraduate perspectives. The remaining responses totals 102 for the Test 2023 cohort and 115 for the Fall 2023 cohort
Participation rates varied across groups, with Test 2023 at an estimated 0.5%, and Fall 2023 at 2.5%. These variations may reflect the differences in email dissemination frequency and engagement levels.
Correlation Analysis: To identify relationships between variables, a correlation matrix was computed using the Pearson correlation coefficient, measuring the linear relationship between two variables. The absolute value of 0.5 was the threshold set to identify pairs of features with high collinearity, ensuring that variables with substantial overlap were flagged for further consideration.
Optional Response Analysis: An inductive coding approach was used to analyze free-response data. Preprocessing included tokenization, lowercasing, and the removal of punctuation and stop words. N-grams (unigrams to quadgrams) were extracted, with low-frequency patterns excluded. Grammatical parsing identified noun and verb phrases, which, alongside n-grams, helped surface recurring themes and patterns.
3 Results
The Fall Undergraduates group is divided into three categories: 28 in Computer Science, 60 in Engineering, and 27 in Other, totaling 115 respondents. In contrast, the Test Computer Science Graduates group consists solely of 102 respondents in Computer Science.
Awareness: A small percentage of respondents in all groups believe GenAI is a short-lived trend, with CompSci participants slightly more likely to agree at 21%. Most respondents view generative AI as a significant advancement, with nearly unanimous agreement among Test participants at 100%, followed by CompSci and Engineer respondents at 96% and 93%, respectively. When asked about the limitations of generative AI’s potential applications, agreement levels vary, with Engineers and Test respondents showing slightly higher agreement at 34% and 37%. Generative AI experience is highest among Test respondents (93%) and lowest among Engineers (78%).
Figure 2: Responses on GenAI awareness.
Preparedness: For the percentage of respondents that believe job skills will change due to GenAI, Fall Undergraduate Computer Science students have the highest percentage at 82%, followed by Test Computer Science Graduates at 77%. Fall Other Majors report 56%, while Fall Engineering students have the lowest percentage at 38%.
Of those that answered “Yes” to job skills changing, there was a follow up question of how soon they believe the changes will occur. Test Computer Science Graduates had the highest expectation of changes occurring the soonest, within 3 years. Fall Undergraduate Computer Science and Other Majors have a more balanced distribution, with nearly half predicting changes within 3 years and the rest expecting shifts within 5 years or longer. The Fall Undergraduate Engineering students mostly expect changes within 5 or 10 years and just 17% expected changes to occur within 3 years.
Figure 3: Respondent counts by degree and cohort.
Figure 4 covers the remaining three questions on GenAI preparedness. Engineering students reported the highest satisfaction with GenAI discussions at 57% and the greatest confidence that professors considered GenAI at 64%. In contrast, Computer Science students had the lowest satisfaction with 38% and the least confidence in faculty engagement at 35%. Despite this, confidence in curriculum preparation was high across all groups, with Engineering leading at 79%. These results suggest stronger GenAI integration in Engineering curricula, while Computer Science students perceive gaps in discussion and faculty consideration.
Figure 4: Responses on preparedness
Concern: The majority view GenAI’s career impact positively, with Test respondents highest at 81% and Engineering, CompSci, and Other lower with 70%, 64%, and 63% respectively. Excitement about GenAI benefits was high across all groups, peaking among Engineering at 77% and Test at 80%. Other respondents notably lower at 67%. Even with optimism and excitement about positive career impact and excitement, concern about GenAI’s impact was also prevalent, with Other disciplines 74% and Engineering 75%, expressing the most caution.
Correlations: Correlations with an absolute value of 0.5 or higher were examined to highlight meaningful patterns in how students view GenAI’s impact. Among Fall undergraduates, excitement about GenAI benefits is moderately correlated (0.51) with perceiving a positive career impact, and satisfaction with GenAI discussions is similarly correlated (0.51) with confidence in professors’ engagement. In computer science, students who believe changes will occur within ten years tend to also view GenAI’s applications as limited, with a correlation of 0.50. Students who feel prepared are more likely to expect positive career outcomes, with a stronger correlation of 0.60. For engineering students, satisfaction with discussions correlates even more strongly with confidence in faculty preparedness, at 0.64. In other disciplines, excitement about GenAI benefits is negatively correlated at -0.57 with the belief that its applications are limited, and positively correlated at 0.76 with a belief in GenAI’s positive career impact. Additionally, satisfaction with discussions is moderately negatively correlated with expecting imminent changes at -0.53 and with anticipating rapid shifts in required skills at -0.55.
Figure 5: Topics raised in free response
Free Response: Optional free response rates were relatively high, with 40–50% of undergraduates and test graduate students responding to the institutional recommendation prompt, and 23–25% responding to the additional thoughts prompt. Due to limited responses by discipline, the Fall undergraduate cohort was analyzed as a whole.
The most common recommendation—mentioned by over 50% of both cohorts—was teaching students how to use GenAI tools. About 33% of respondents also suggested integrating GenAI into existing curricula or adding new courses. Ethics and career implications of GenAI appeared more frequently in undergraduate responses, with ethics mentioned 15% more and career impact 5% more than in graduate responses.
Emotional tone varied between cohorts. Over half of undergraduates expressed negative emotions, 29% more than the test graduate group, while test graduates were 20% more likely to express positivity. Additional thoughts covered a broader range of themes. Test graduates frequently emphasized critical thinking, whereas undergraduates more often expressed uncertainty.
These open-ended responses provide rich context to the survey data and are further explored in the Discussion. While the first prompt focused on GenAI’s utility, the second invited any remaining reflections from participants.
4 Discussion
The results have presented unique and shared perspectives across undergraduate programs, revealing areas of improvement. Thanks to the participants in this study and their contributions to the free response, there is a story that has begun to emerge surrounding the needs of the GenAI Generation as they learn and grow into their future careers.
Awareness: the uncertainty in early adopters
While overall awareness of GenAI is high, this awareness does not translate uniformly into understanding or confidence. Non-STEM students exhibited the strongest correlation between excitement about GenAI and its perceived career impact, suggesting an optimistic yet potentially less critically informed perspective. In contrast, engineering students had the highest skepticism about GenAI’s practical applications, while computer science students were most likely to view GenAI as a short-lived trend. These viewpoints may reflect the varied depth of engagement with AI technologies in their respective fields, with technical students balancing enthusiasm with caution about GenAI’s limitations and ethical implications.
These nuances are echoed in free responses, where students shared diverse sentiments:
“because aerospace engineering is pretty specialized and oftentimes requires very difficult/complex work, i am ultimately unsure how genai would be used in my field, if at all. however, i trust it can be a valuable tool in the future. i don’t feel like genai has any negative impacts on my future.” — Undergraduate Engineer
“i think it has limited application in anything with more advanced logic or understanding.” — Undergraduate Engineer
“i think it has limited potential because genai currently cannot distinct whether the information it gives is true or not.” — Undergraduate Engineer
“it’s very uncertain and that’s what makes me slightly afraid. i also don’t understand it very well, which i should work on.” — Undergraduate Computer Science
“as far as i can tell, it’s a fad for almost all possible proposed use cases. until stronger evidence comes to light for its all-encompassing power, it should be ignored in higher education.” — Undergraduate Engineer
As seen in some of these responses, there is skepticism of GenAI achieving beyond what the current potential shows, lending to an experience of technology moving faster beyond students’ grasp. Many students rely on self-directed learning, coursework, and personal experiences to form their perspectives on GenAI. This gap in structured AI education leads to a divergence in awareness, confidence, and preparedness, with some students feeling optimistic and others expressing deep concern and uncertainty.
In the optional free responses, 50% of the students expressed negative emotions, including frustration, fear, and worry, while 20% conveyed uncertainty, with overlapping sentiments highlighting the complex emotional landscape students navigate:
“i work in security. i am not excited about what’s in store.” — Undergraduate Other
“i’m scared it’s going to make my career not what i want it anymore.” — Undergraduate Other
These comments reflect fear and uncertainty about GenAI’s impact, amplified by a lack of classroom discussion. Without institutional guidance, students are left to navigate these changes on their own during critical educational years. Some expressed dissatisfaction with the current state of awareness in their programs overall:
“even addressing its existence in class would be a step forward.” — Undergraduate Engineer
For a transformative tool like GenAI, students are actively seeking guidance to understand its potential, limits, and ethical implications.
Preparedness: addressing the elephant in the classroom
While students recognize GenAI’s growing influence on their careers, many remain unsure whether their education is preparing them for the changes ahead. Unlike future students who may benefit from GenAI-integrated curricula, the current generation often relies on self-directed strategies to prepare for a GenAI-driven workforce.
This preparedness gap is reflected in how different groups perceive GenAI’s impact on job skills. Among computer science students, there is a clear pattern: those who feel prepared are most likely to expect a positive career impact.
In contrast, only 56% of ”Other” students and 38% of engineering students anticipate such changes, with many engineering students expecting a slower pace—highlighting a lack of urgency in their field. These differences suggest the need for tailored strategies that align with each group’s career context.
Despite these varied views, all cohorts reported relatively high confidence in their curricula—possibly reflecting trust in their degree or Institutions’ reputation. Still, engineering students stand out for their skepticism: they are the least likely to expect GenAI-driven skill changes, report the lowest satisfaction with classroom discussions, and express the least confidence in faculty engagement. This gap signals a need for deeper integration and clearer guidance.
Correlations support these patterns. In engineering, classroom discussions on GenAI are linked to higher confidence in professors, suggesting that open engagement builds trust and a sense of readiness. In computer science, skepticism often aligns with delayed expectations, while students who feel prepared tend to view GenAI as a career asset. In the ”Other” cohort, satisfaction with discussions correlates with the belief that GenAI will impact job skills, suggesting confidence in the status quo or comfort with limited exposure.
The free response data adds depth to these findings. Despite overall confidence in their curriculum, over half of students recommended more courses on GenAI and its practical use. Many also called for stronger emphasis on ”fundamentals,” ”critical thinking,” and ethical considerations:
“I think focusing on the idea of working alongside genai on how to take advantage of it to produce better, faster results, but acknowledge its limitations […] and discussing ethical and other ramifications of using ai in certain situations. — Undergraduate Other
“it is a ”tool” that can be used to aid students’ work when used responsibly, and can be integrated in courses like other software tools, such as matlab, cad softwares, etc. but should not replace the fundamentals of engineering that is already being taught, such as hand calculations and understanding the theory.” — Undergraduate Engineer
“strong implementation in to curriculum and learning tools to learn to work with ai and not treat it as a monster.” — Undergraduate Engineer
“[…] the more adapted we are to it and prepared for its growth/advancement, the better off we’ll be to use it as a tool, rather than as a toy. — Undergraduate Other
While many responses showed a proactive approach to GenAI integration, others expressed skepticism or disappointment:
“[…]i’m disappointed that tech won’t prepare me for the near advent of those tools during my classes.” — Undergraduate Engineer
“as far as i can tell, it’s a fad for almost all possible proposed use cases. until stronger evidence comes to light for its all-encompassing power, it should be ignored in higher education.” — Undergraduate Engineer
“show the downsides and flaws of things like gpt-4, discuss environmental impacts, and don’t use it. genai is built on stolen data sets and harms artists. stop endorsing its use. — Undergraduate Other
The findings reveal a misalignment between perceived career impact and readiness. As GenAI begins to enter technical curricula, students in other disciplines risk being left behind. Addressing this gap will require expanding GenAI education across fields, embedding GenAI literacy into non-STEM programs, and equipping faculty to lead informed, inclusive discussions. These efforts are essential to preparing a more equitable and capable generation for a GenAI-driven future.
Concern: learning at what cost?
Student concerns about GenAI range from ethics and job security to institutional and faculty preparedness. While excitement is widespread, it is tempered by deeper anxieties, reflecting the complexity of the GenAI Generation’s experience. The Test 2023 Computer Science Graduate cohort reported the lowest concern (65%) and the highest levels of excitement (80%) and belief in GenAI’s positive career impact (81%). This likely reflects the integration of GenAI into their curriculum and their confidence using such tools.
“[…] genai may be able to do the bulk of the work when designing something functional and usable, so the focus can now be on designing technology that people like and that benefits them. i think genai can help me do what i like, which would benefit people and society.“ — Graduate Computer Science
“im excited for the changes and technological advances it will bring.“ — Graduate Computer Science
In contrast, positive outlook drops to 64% among undergraduate computer science students, reinforcing how exposure shapes perception.
Among undergraduates, belief in a positive career impact ranges from 63% to 70%, with engineering students on the higher end. Excitement is highest among computer science (75%) and engineering students (77%), while the “Other” cohort remains steady at 67%. Even when students are unsure about GenAI’s direct impact on their careers, many still express enthusiasm about its broader potential. Some specific positive takes included:
“i think less people should have to work for as long and more leisure and volunteering time will be helpful for mental health and balance.” — Undergraduate Engineer
“GenAI is the beginning of the 3rd “industrial” revolution. My life has changed a lot from ChatGPT.” — Undergraduate Other
“i think it is important to emphasize both positives and negatives of genai whenever mentioned. ” — Undergraduate Computer Science
Free responses reflect strong concern, especially in the “additional thoughts” section. Over 50% conveyed negative emotions, with recurring themes of ethics, job replacement, and uncertainty—often expressed in long, passionate comments.
“i think it will actively damage the quality of work and data produced[…]” — Undergraduate Engineer
“to be honest, i really do not care. i feel like genai is going to obliterate most jobs. it almost seems inevitable.[…]the only thing holding me from hopelessness is (blind) faith that my role as a researcher will not be rendered obsolete.” — Undergraduate Computer Science
“i think we are on a precipice of either ruin or greatness. regardless of how we use it, genai will become a powerful tool and have an impact greater than any singular tool since humans gained control of fire.[…] ”— Undergraduate Other
Some students called for regulation:
“[…]we must ensure that there are precautions before the ai reaches a level where its misuse could pose a legitimate harm towards society and overall human advancement.” — Undergraduate Other
Others expressed concern about government use:
“i think the governments will start using genai to produce propoganda in the coming years (if they haven’t already) to manipulate public opinion. or maybe just politicians with bot accounts blindly supporting them and arguing with dissenters on social media.” — Undergraduate Engineer
And some of the more succinct, poignant responses:
“goodbye middle class, we doomed.” — Undergraduate Computer Science
“i’m scared it’s going to make my career not what i want it anymore.” — Undergraduate Other
Student responses reveal strong concerns about the ethical, economic, and societal impacts of GenAI, alongside a desire for greater guidance and critical engagement. Institutions must actively support both students and faculty as they navigate this rapidly evolving landscape.
5 Conclusion
The GenAI Generation, defined as students in their formative educational years during the rise of GenAI, face a rapidly evolving landscape shaped by both opportunity and uncertainty. While many are excited about GenAI’s potential, they also express concerns around preparedness, ethics, and long-term implications. Findings show that students with greater exposure to GenAI, particularly graduate computer science students, report higher confidence and optimism about their future careers. However, such structured exposure is not consistent across disciplines, revealing an institutional gap that must be addressed to ensure equitable support for all students navigating this technological shift.
The GenAI Generation is distinct in its developmental context as students are not only adapting to new technologies but doing so without established norms or clear guidelines. They are the early adopters of tools like ChatGPT, learning to balance traditional skills with new competencies required to interact with GenAI effectively. Their adaptability is frequently tested while grappling with ethical questions, workforce changes, and broader societal implications.
Students want more than just exposure: they seek guidance on using GenAI tools responsibly, critical thinking around AI-generated outputs, and a robust understanding of ethical considerations. The GenAI Generation is asking for a holistic approach that not only enhances technical skills but also nurtures critical engagement and ethical reasoning. It is critical to closely monitor and derive clarity on the potential impacts of these tools, developing a deeper understanding of cognition and ethical implications.
The GenAI Generation is at a pivotal moment. As the first cohort to learn and grow in a GenAI-centric world, their experiences will set the tone for future generations. The data highlights the need for GenAI literacy efforts that span disciplines and prepare both students and faculty to engage with these tools in meaningful ways. The voices of these students are clear: they want to be part of a future where GenAI is a tool for growth, not a source of uncertainty, and where their education equips them to use GenAI as a force for positive change.
Acknowledgements
We would like to thank the students who took the time to respond to our survey, Dr. Joseph Ludlum and Dr. Loraine Phillips from the Office of Academic Effectiveness at Georgia Tech for their support in distributing the Fall 2023 cohort survey, and Dr. Thomas Ploetz for the continued encouragement throughout the process.

### v2

@@
 Abstract
-Generative AI (GenAI) is revolutionizing education and workforce development, profoundly shaping how students learn, engage, and prepare for their future.
-Outpacing the development of uniform policies and structures, GenAI has heralded a unique era and given rise to the GenAI Generation: a cohort of students whose education has been increasingly shaped by the opportunities and challenges GenAI presents during its widespread adoption within society.
-This study examines our students’ perceptions of GenAI through a concise survey with optional open-ended questions, focusing on their awareness, preparedness, and concerns.
-Evaluation of more than 250 responses with more than 40% providing detailed qualitative feedback reveals a core dual sentiment: while most students express enthusiasm for GenAI, an even greater proportion voice a spectrum of concerns about ethics, job displacement, and the adequacy of educational structures given the highly transformative technology.
-These findings offer critical insights into how students view the potential and pitfalls of GenAI for future career impacts, with accompanying recommendations to guide educational institutions in navigating a future driven by GenAI.
+Generative Artificial Intelligence (GenAI) is revolutionizing education and workforce development, profoundly shaping how students learn, engage, and prepare for their future.
+Outpacing the development of uniform policies and structures, GenAI has heralded a unique era and given rise to the GenAI Generation.
+We define the GenAI Generation as a cohort of students whose education has been increasingly shaped by the opportunities and challenges GenAI presents during its widespread adoption within society.
+This study examines students’ perceptions of GenAI through a concise survey with optional open-ended questions, focusing on their awareness, preparedness, and concerns.
+Notably, readiness appears increasingly tied to exposure to GenAI through one’s coursework.
+Students with greater curricular exposure to GenAI tend to feel more prepared, while those without it more often express vulnerability and uncertainty, highlighting a new and growing divide in readiness that goes beyond traditional disciplinary boundaries.
+Evaluation of more than 250 responses, with over 40% providing detailed qualitative feedback, reveals a core dual sentiment: while most students express enthusiasm for GenAI, an even greater proportion voice a spectrum of concerns about ethics, job displacement, and the adequacy of educational structures given the highly transformative technology.
+These findings offer critical insights into how students view the potential and pitfalls of GenAI for future career impacts.
+The challenge ahead involves implementing associated recommendations for educational institutions, moving beyond the baseline of access toward more informed guidance on the use of these tools, while preserving critical thinking, ethical reasoning, and adaptive learning.
+Keywords: Generative Artificial Intelligence; GenAI Generation; Student Perceptions; Higher Education; Career Readiness; Cognitive Development; Adaptability; Problem-Solving; Critical Thinking; Ethics in AI
 1 Introduction
-The rapid advancement of artificial intelligence (AI), particularly generative AI (GenAI), is reshaping the educational landscape.
+The significant advancement of Artificial Intelligence (AI), particularly Generative AI (GenAI), is reshaping the educational landscape.
+GenAI is a type of AI trained on large datasets to generate new content such as text, images, audio, or video in response to user prompts, setting it apart from predictive models by its ability to create [22, 28, 20, 21].
 As a tool capable of autonomously producing human-like content, GenAI offers both opportunities and uncertainties for education.
 It can enhance creativity, streamline workflows, and support personalized learning, yet its broader impact remains unclear.
-As GenAI becomes more embedded in academic and professional contexts, understanding how students perceive its influence on future readiness is essential.
+As GenAI continues to evolve exponentially and integrate within academic contexts, there is a pressing need to understand how it is shaping the developmental experiences of students in their formative academic years.
+Unlike previous waves of technology’s impact on education, GenAI’s pace and scale introduce a layer of uncertainty around its long-term effects on cognition, critical thinking, and engagement.
+By honing in on student perspectives, this research provides early insight into how GenAI is influencing not only perceived career readiness but also the cognitive processes underpinning learning and decision-making.
 This study presents the development, dissemination, and evaluation of a survey examining student perspectives on GenAI through three core lenses: awareness, preparedness, and concern.
-By analyzing responses, this research provides actionable insights into how the GenAI Generation, students most affected by GenAI’s rapid integration, understand its role in shaping their academic and career paths.
-Defining the GenAI Generation
+By analyzing responses, it provides actionable insights into how the GenAI Generation—students most directly impacted by the rise of GenAI—understand its role in shaping both their academic experiences and future trajectories.
+The GenAI Generation
 The GenAI Revolution represents a paradigm shift in technology adoption, enhancing efficiency, creativity, and performance while introducing challenges such as ethical concerns and the need for responsible stewardship [11, 5, 16, 4].
 As it relates to education, this revolution not only emphasizes a transformation in learning but also highlights the challenges students and teachers face during a phase of unclear uniform guidance [4, 5].
-This study introduces the concept of the GenAI Generation to describe individuals most directly impacted by the Generative AI Revolution.
+This study introduces the GenAI Generation.
 The GenAI Generation is defined as the cohort of individuals actively undergoing developmental and educational growth/attainment during the rise of Generative AI, starting with the release of ChatGPT in November 2022.
@@
   • Navigating Uncertainty: Early adopters, engaging with experimental, unregulated technologies without widely established policies.
-  • High Adaptability: Balancing both the traditional and new competencies required to effectively interact and engage with GenAI in a rapidly changing environment.
+  • High Adaptability: Balancing both the traditional and new competencies required to effectively interact and engage with GenAI in a highly dynamic environment.
   • Long-Term Impact Unknown: Using these tools during formative years without a clear understanding of potential long-term cognitive or social consequences.
@@
 Contributions of This Study
-This study contributes to the ongoing dialogue about GenAI’s role in shaping students’ academic and professional paths.
-Centered on awareness, preparedness, and concern, the survey offers a structured view of the GenAI Generation, a group navigating rapid technological change during key years of education.
-Focusing on higher education, the analysis shows how students weigh GenAI’s potential to support creativity and critical thinking against concerns about overreliance, ethical risks, and reduced engagement.
-These insights emphasize the need for thoughtful strategies that support ethical and effective use of GenAI in education.
+This study offers timely insights into how students in their formative educational years are experiencing the rapid rise of Generative AI.
+It introduces the term “GenAI Generation” to describe students navigating both academic development and technological transformation without consistent institutional guidance or established norms.
+Drawing from a combination of structured survey data and rich qualitative responses, the study reveals meaningful patterns in students’ awareness, preparedness, and concern regarding GenAI.
+Across majors, students expressed a desire for more intentional support, not just access to GenAI tools, but structured opportunities to learn how to use them responsibly.
+While formal curriculum integration remains inconsistent, students who encounter GenAI through coursework or self-directed exploration are more likely to articulate confidence and optimism about its role in their future.
+From enthusiasm to anxiety, the findings illuminate the psychosocial and pedagogical dimensions of GenAI’s presence in higher education.
+Students are not only engaging with a new technology, they are forming perceptions of their own readiness, trust in institutions, and ethical responsibility in real time.
+These perspectives carry clear implications for curriculum design, faculty preparation, and equitable access to digital literacy.
+This work contributes to an emerging understanding of what students need to thrive in a GenAI-influenced world: not only technical fluency, but critical engagement, ethical grounding, and institutional clarity.
+The study underscores the importance of proactive educational responses that extend beyond STEM fields to support students across all disciplines in this pivotal moment.
 2 Methods
@@
 Awareness, preparedness, and concern guided the development of eight main questions, three demographic items, and two open-ended prompts.
+This study received Institutional Review Board (IRB) approval from the [Institution Name Redacted for Review], and all participants provided informed consent prior to beginning the survey.
 Designed for completion in under five minutes [13], the survey began with an IRB-required disclosure and a screening question, followed by items progressing from general to sensitive topics [24].
@@
 While omitting a neutral option is debated [14], this was an intentional design choice to encourage engagement and reduce bias [15].
-The survey was administered across two cohorts, with full question text available in the supplemental materials.
-The Test 2023 survey was distributed to Computer Science Master’s students via OMSCS (12,000 members) and MSCS (14,000 members) listservs.
+The survey was administered across two [Institution Name Redacted for Review] student cohorts, with full question text available in the supplemental materials.
+The first cohort, Fall 2023 Computer Science Graduate students, survey was distributed to Online Computer Science Master’s (OMSCS) students, approximately 12,000 members [CITE REDACTED FOR REVIEW], and campus Computer Science Master’s (MSCS), approximately 14,500 members [CITE REDACTED FOR REVIEW], via listservs.
 Invitations were sent on September 28 and 29, 2023, with responses collected through November 1.
-The Fall 2023 survey was conducted with support from the Office of Academic Effectiveness, which used stratified random sampling to select 5,655 undergraduates representative of the broader student population.
+The second cohort, Fall 2023 Undergraduate students, survey was conducted with support from the [Institution Name Redacted for Review] Office of Academic Effectiveness (OAE), which used stratified random sampling to select 5,655 undergraduates representative of the broader student population.
 Sampling quotas ensured balance by gender, race, citizenship, and class standing, with a 2% margin of error.
+The demographic breakdown of the sample can be found in the supplemental materials.
 The survey was open from November 2–23, 2023, with email reminders sent on November 2, 15, and 22.
-Methods of Analysis
-Survey responses were analyzed using correlation, clustering, and free-response evaluation to capture both quantitative and qualitative insights.
-These methods supported a systematic exploration of trends and perceptions across diverse student groups.
+Analysis Approach
+Descriptive statistics and frequency distributions were used to compare response patterns across student subgroups, while qualitative evaluation of open-ended responses enabled a systematic exploration of emerging themes and perceptions.
 Data Processing: To ensure data integrity, only completed responses containing degree type information were included in the analysis.
-Degrees were categorized into three groups: Computer Science, Engineering, and Other, based on thematic alignment with disciplinary focus and methodologies.
-Multidisciplinary fields, such as Robotics and Computational Science and Engineering, were classified under Engineering and Computer Science, respectively, to maintain consistency.
-This classification enabled cross-disciplinary comparisons and analysis of AI’s impact on skills by degree category and education level.
-For the Fall 2023 cohort, two non-undergraduate entries, likely from dual-enrolled students, were excluded to maintain a focus on undergraduate perspectives.
-The remaining responses totals 102 for the Test 2023 cohort and 115 for the Fall 2023 cohort
-Participation rates varied across groups, with Test 2023 at an estimated 0.5%, and Fall 2023 at 2.5%.
-These variations may reflect the differences in email dissemination frequency and engagement levels.
-Correlation Analysis: To identify relationships between variables, a correlation matrix was computed using the Pearson correlation coefficient, measuring the linear relationship between two variables.
-The absolute value of 0.5 was the threshold set to identify pairs of features with high collinearity, ensuring that variables with substantial overlap were flagged for further consideration.
-Optional Response Analysis: An inductive coding approach was used to analyze free-response data.
+Where the first cohort of graduate students comprised soley of computer science majors, the undergraduate cohort had many degrees.
+Rather than broadly distinguishing between STEM and non-STEM, degrees were categorized into Computer Science, Engineering, and Other, based on disciplinary focus and methodological orientation.
+Table 1 provides a Major-Category mapping for reference.
+This grouping structure reflects both practical sample sizes (Other, 28 Computer Science, and 60 Engineering) and a meaningful distinction in how students may engage with GenAI in academic contexts.
+This also enabled a comparison between computer science graduate and undergraduate perspectives.
+Grouping a diverse range of degrees—such as Biology, Business Administration, Literature, and Public Policy—under the single category “Other” introduces a potential limitation, as it may obscure important disciplinary differences in students’ exposure to or perspectives on GenAI.
+While specific exposure to AI tools and curricula was not measured directly, these categories serve as a reasonable proxy for varying degrees of disciplinary integration with GenAI-related content, allowing exploratory comparisons across educational domains.
+Table 1: Undergraduate Sample and Response Rates by Majors and Category
+Major* Category Sample Freq. Response Freq. Sample % Response %
+Applied Physics Other 12 2 0.2 1.7
+Biology Other 162 3 2.9 2.6
+Business Administration Other 440 3 7.8 2.6
+Computational Media Other 98 1 1.7 0.9
+Earth & Atmospheric Sciences Other 13 1 0.2 0.9
+Econ & Int’l Affairs Other 15 1 0.3 0.9
+Economics Other 49 2 0.9 1.7
+History, Technology, & Society Other 18 2 0.3 1.7
+Industrial Design Other 88 2 1.6 1.7
+Int’l Affairs & Mod Lang Other 18 2 0.3 1.7
+Lit., Media, & Communication Other 59 2 1.0 1.7
+Mathematics Other 77 3 1.4 2.6
+Neuroscience Other 146 3 2.6 2.6
+Physics Other 68 5 1.2 4.3
+Psychology Other 66 5 1.2 4.3
+Public Policy Other 37 1 0.7 0.9
+Aerospace Engineering Engineer 349 8 6.2 7.0
+Biomedical Engineering Engineer 399 6 7.1 5.2
+Chemical and Biomolecular Eng Engineer 187 5 3.3 4.3
+Computer Engineering Engineer 187 2 3.3 1.7
+Electrical Engineering Engineer 187 3 3.3 2.6
+Environmental Engineering Engineer 58 3 1.0 2.6
+Industrial Engineering Engineer 344 5 6.1 4.3
+Materials Science & Engr Engineer 67 3 1.2 2.6
+Mechanical Engineering Engineer 582 12 10.3 10.4
+Nuclear & Radiological Engr Engineer 32 2 0.6 1.7
+Computer Science Computer Science 1419 28 25.1 24.3
+Applied Lang/Intercultural St No Response 4 0 0.1 0.0
+Architecture No Response 76 0 1.3 0.0
+Atmospheric & Oceanic Sciences No Response 3 0 0.1 0.0
+Biochemistry No Response 88 0 1.6 0.0
+Building Construction No Response 13 0 0.2 0.0
+Chemistry No Response 40 0 0.7 0.0
+Environmental Science No Response 10 0 0.2 0.0
+Global Econ/Mod Lang No Response 5 0 0.1 0.0
+International Affairs No Response 22 0 0.4 0.0
+Music Technology No Response 23 0 0.4 0.0
+Solid Earth/Planetary Sciences No Response 2 0 0.0 0.0
+Special/Non-Degree No Response 64 0 1.1 0.0
+Civil Engineering No Response 128 0 2.3 0.0
+*For reference; sample was not balanced on this attribute. Sample breakdowns by additional attributes can be found within the supplemental material.
+For the undergraduate cohort, two non-undergraduate entries, likely from dual-enrolled students, were excluded to maintain a focus on undergraduate perspectives.
+The remaining responses totals 102 for the graduate cohort and 115 for the undergraduate cohort.
+Table 2 shows this process
+Table 2: Table showing data cleaning process
+Survey Population Total Started Total Incomplete Total Dropped Reason for Drop Remaining Completed Total
+Computer Science Graduates 2023 26,000** 122 19 1 Response was “prefer not to answer” for degree 102
+Undergraduates 2023 5,655 143 26 2 Did not have “undergraduate” for degree type 115
+** Approximate total population estimate for Computer Science Graduate 2023 cohort.
+Qualitative Free-Response Analysis: An inductive coding approach was used to analyze free-response data.
 Preprocessing included tokenization, lowercasing, and the removal of punctuation and stop words.
@@
 3 Results
-The Fall Undergraduates group is divided into three categories: 28 in Computer Science, 60 in Engineering, and 27 in Other, totaling 115 respondents.
-In contrast, the Test Computer Science Graduates group consists solely of 102 respondents in Computer Science.
-Awareness: A small percentage of respondents in all groups believe GenAI is a short-lived trend, with CompSci participants slightly more likely to agree at 21%.
-Most respondents view generative AI as a significant advancement, with nearly unanimous agreement among Test participants at 100%, followed by CompSci and Engineer respondents at 96% and 93%, respectively.
-When asked about the limitations of generative AI’s potential applications, agreement levels vary, with Engineers and Test respondents showing slightly higher agreement at 34% and 37%.
-Generative AI experience is highest among Test respondents (93%) and lowest among Engineers (78%).
+The undergraduate cohort included 115 respondents across three categories: 28 in Computer Science, 60 in Engineering, and 27 in Other disciplines.
+The graduate cohort response count was 102 and comprised solely of computer science majors.
+The full breakdown between sample to response counts can be seen in Table 1.
+Figure 2: Response counts by degree and level
+Awareness: A small percentage of respondents in all groups believe GenAI is a short-lived trend, seen in Figure 3, with undergraduate computer science students slightly more likely to agree at 21%.
+Most respondents view generative AI as a significant advancement, with nearly unanimous agreement among the graduate cohort at 100%, followed by undergraduate computer science and engineers at 96% and 93%, respectively.
+When asked about the limitations of generative AI’s potential applications, agreement levels vary, with engineers and graduate respondents showing slightly higher agreement at 34% and 37%.
+Generative AI experience is highest among Graduate respondents (93%) and lowest among Engineers (78%).
 Figure 3: Responses on GenAI awareness.
-Preparedness: For the percentage of respondents that believe job skills will change due to GenAI, Fall Undergraduate Computer Science students have the highest percentage at 82%, followed by Test Computer Science Graduates at 77%.
-Fall Other Majors report 56%, while Fall Engineering students have the lowest percentage at 38%.
+Preparedness: For the percentage of respondents that believe job skills will change due to GenAI, Figure 4 shows undergraduate computer science students have the highest percentage at 82%, followed by computer science graduates at 77%.
+Other majors report 56%, while engineering students have the lowest percentage at 38%.
+Figure 4: Responses on if job skills will change due to GenAI.
 Of those that answered “Yes” to job skills changing, there was a follow up question of how soon they believe the changes will occur.
-Test Computer Science Graduates had the highest expectation of changes occurring the soonest, within 3 years.
-Fall Undergraduate Computer Science and Other Majors have a more balanced distribution, with nearly half predicting changes within 3 years and the rest expecting shifts within 5 years or longer.
+Figure 5 shows computer science graduates had the highest expectation of changes occurring the soonest.
+Undergraduate Computer Science and Other Majors have a more balanced distribution, with nearly half predicting changes within 3 years and the rest expecting shifts within 5 years or longer.
 The Fall Undergraduate Engineering students mostly expect changes within 5 or 10 years and just 17% expected changes to occur within 3 years.
-Figure 3: Respondent counts by degree and cohort.
 Figure 6 covers the remaining three questions on GenAI preparedness.
@@
 These results suggest stronger GenAI integration in Engineering curricula, while Computer Science students perceive gaps in discussion and faculty consideration.
+Figure 5: Respondent counts by degree and cohort.
 Figure 6: Responses on preparedness
-Concern: The majority view GenAI’s career impact positively, with Test respondents highest at 81% and Engineering, CompSci, and Other lower with 70%, 64%, and 63% respectively.
+Concern: The majority view GenAI’s career impact positively, seen in Figure 7, with computer science graduates highest at 81% and undergraduate Engineers, Computer Science, and Other lower with 70%, 64%, and 63% respectively.
 Excitement about GenAI benefits was high across all groups, peaking among Engineering at 77% and Test at 80%.
 Other respondents notably lower at 67%.
-Even with optimism and excitement about positive career impact and excitement, concern about GenAI’s impact was also prevalent, with Other disciplines 74% and Engineering 75%, expressing the most caution.
-Correlations: Correlations with an absolute value of 0.5 or higher were examined to highlight meaningful patterns in how students view GenAI’s impact.
-Among Fall undergraduates, excitement about GenAI benefits is moderately correlated (0.51) with perceiving a positive career impact, and satisfaction with GenAI discussions is similarly correlated (0.51) with confidence in professors’ engagement.
-In computer science, students who believe changes will occur within ten years tend to also view GenAI’s applications as limited, with a correlation of 0.50.
-Students who feel prepared are more likely to expect positive career outcomes, with a stronger correlation of 0.60.
-For engineering students, satisfaction with discussions correlates even more strongly with confidence in faculty preparedness, at 0.64.
-In other disciplines, excitement about GenAI benefits is negatively correlated at -0.57 with the belief that its applications are limited, and positively correlated at 0.76 with a belief in GenAI’s positive career impact.
-Additionally, satisfaction with discussions is moderately negatively correlated with expecting imminent changes at -0.53 and with anticipating rapid shifts in required skills at -0.55.
-Figure 5: Topics raised in free response
-Free Response: Optional free response rates were relatively high, with 40–50% of undergraduates and test graduate students responding to the institutional recommendation prompt, and 23–25% responding to the additional thoughts prompt.
-Due to limited responses by discipline, the Fall undergraduate cohort was analyzed as a whole.
-The most common recommendation—mentioned by over 50% of both cohorts—was teaching students how to use GenAI tools.
+Even with excitement and optimism about positive career impacts, concern was even greater across all undergraduate degree categories.
+While less, the computer science graduate cohort still showed concern with 65% concerned with GenAI’s.
+Figure 7: Responses on concern
+Free Response: In addition to the structured survey responses, a notable portion of participants contributed to the optional open-ended questions.
+These qualitative insights deepen understanding of the perspectives behind the patterns and are summarized below.
+40–50% of undergraduates and graduate students responded to the institutional recommendation prompt, and 23–25% provided input on the additional thoughts prompt.
+Participation rates for the free response can be seen in Table 3, showing non-Computer Science students having the highest participation.
+The Fall undergraduate free responses were analyzed as a whole.
+Table 3: Optional Free Response Participation by Question & Degree Category
+Category Institutional Recommendation Total Additional Thoughts Total Survey Response Total
+Undergraduate Other 17 11 38
+Undergraduate Engineer 24 11 49
+Undergraduate Computer Science 6 5 29
+Graduate Computer Science 52 27 102
+Figure 8: Topics raised in free response for institutional recommendations
+The most common recommendation, mentioned by over 50% for both cohorts, was teaching students how to use GenAI tools (Figure 8.
 About 33% of respondents also suggested integrating GenAI into existing curricula or adding new courses.
 Ethics and career implications of GenAI appeared more frequently in undergraduate responses, with ethics mentioned 15% more and career impact 5% more than in graduate responses.
-Emotional tone varied between cohorts.
-Over half of undergraduates expressed negative emotions, 29% more than the test graduate group, while test graduates were 20% more likely to express positivity.
-Additional thoughts covered a broader range of themes.
-Test graduates frequently emphasized critical thinking, whereas undergraduates more often expressed uncertainty.
+There were fewer who responded to the optional ”additional thoughts” question, those who did showed clear emotional tone(Figure 9).
+Over half of undergraduates expressed negative emotions, 29% more than the graduate cohort, while graduates were 20% more likely to express positivity.
+Additional thoughts covered a broader range of themes(Figure 10).
+The computer science graduates frequently emphasized critical thinking, whereas undergraduates more often expressed uncertainty.
+Figure 9: Percentage of free responses with either positive or negative emotions conveyed
+Figure 10: Topics raised in free response additional thoughts
 These open-ended responses provide rich context to the survey data and are further explored in the Discussion.
@@
 Thanks to the participants in this study and their contributions to the free response, there is a story that has begun to emerge surrounding the needs of the GenAI Generation as they learn and grow into their future careers.
+Although surveys were distributed uniformly to ensure equal access across cohorts, the potential for selection bias and low response frequencies remains a limitation.
+Response counts were modest compared to sample size, and some subgroups—particularly within non-CS or less tech-aligned majors—were underrepresented.
+As a technology-focused institution, [Institution Name Redacted for Review]’s student population may also have greater exposure to emerging technologies, including GenAI, than students at other institutions.
+Even with these constraints, the findings offer important early insight into how GenAI is shaping student experiences, perceptions, and expectations during a critical period of educational and technological transformation.
 Awareness: the uncertainty in early adopters
-While overall awareness of GenAI is high, this awareness does not translate uniformly into understanding or confidence.
-Non-STEM students exhibited the strongest correlation between excitement about GenAI and its perceived career impact, suggesting an optimistic yet potentially less critically informed perspective.
-In contrast, engineering students had the highest skepticism about GenAI’s practical applications, while computer science students were most likely to view GenAI as a short-lived trend.
-These viewpoints may reflect the varied depth of engagement with AI technologies in their respective fields, with technical students balancing enthusiasm with caution about GenAI’s limitations and ethical implications.
+While overall awareness of GenAI was high across all groups, levels of understanding and confidence varied.
+Computer science undergraduates were the most likely to consider GenAI a short-lived trend (21%), whereas engineering and “Other” majors reported lower agreement (10–11%).
+Despite this skepticism, nearly all respondents across disciplines viewed GenAI as a meaningful advancement, with agreement ranging from 89% to 100%.
+Experience using GenAI tools also varied, with graduate students reporting the highest exposure (93%) and engineering undergraduates the lowest (78%).
+Responses regarding the limitations of GenAI’s potential were more common among engineering and graduate students.
+For the graduate cohort comprising of computer science, this could suggest increased exposure may come with greater critical awareness of its boundaries.
+For the undergraduate engineering students, skepticism may come from the belief that GenAI’s application is limited specifically for their degree.
 These nuances are echoed in free responses, where students shared diverse sentiments:
@@
 “It’s very uncertain and that’s what makes me slightly afraid. I also don’t understand it very well, which I should work on.” — Undergraduate Computer Science
-“as far as i can tell, it’s a fad for almost all possible proposed use cases.
-until stronger evidence comes to light for its all-encompassing power, it should be ignored in higher education.” — Undergraduate Engineer
 As seen in some of these responses, there is skepticism of GenAI achieving beyond what the current potential shows, lending to an experience of technology moving faster beyond students’ grasp.
@@
 “Even addressing its existence in class would be a step forward.” — Undergraduate Engineer
+These emotional responses and findings reveal a misalignment between perceived career impact and readiness and signal a broader psychosocial landscape in flux where students must construct academic identity and career expectations while navigating an unregulated and evolving technology.
 For a transformative tool like GenAI, students are actively seeking guidance to understand its potential, limits, and ethical implications.
@@
 While students recognize GenAI’s growing influence on their careers, many remain unsure whether their education is preparing them for the changes ahead.
-Unlike future students who may benefit from GenAI-integrated curricula, the current generation often relies on self-directed strategies to prepare for a GenAI-driven workforce.
+Unlike future students who may benefit from GenAI-integrated curricula, the GenAI Generation often relies on self-directed strategies to prepare for a GenAI-driven workforce.
 This preparedness gap is reflected in how different groups perceive GenAI’s impact on job skills.
-Among computer science students, there is a clear pattern: those who feel prepared are most likely to expect a positive career impact.
-In contrast, only 56% of ”Other” students and 38% of engineering students anticipate such changes, with many engineering students expecting a slower pace—highlighting a lack of urgency in their field.
+There is a clear pattern that begins to emerge: those who feel prepared in their degree are most likely to expect a positive career impact.
+Where computer science 77% of graduates and 82% undergraduates anticipate job skill changes, only 56% and 38% of undergaduate other and engineering students anticipate such changes, with many engineering students expecting a slower pace—highlighting a lack of urgency in their field.
 These differences suggest the need for tailored strategies that align with each group’s career context.
@@
 This gap signals a need for deeper integration and clearer guidance.
-Correlations support these patterns.
-In engineering, classroom discussions on GenAI are linked to higher confidence in professors, suggesting that open engagement builds trust and a sense of readiness.
-In computer science, skepticism often aligns with delayed expectations, while students who feel prepared tend to view GenAI as a career asset.
-In the ”Other” cohort, satisfaction with discussions correlates with the belief that GenAI will impact job skills, suggesting confidence in the status quo or comfort with limited exposure.
 The free response data adds depth to these findings.
-Despite overall confidence in their curriculum, over half of students recommended more courses on GenAI and its practical use.
-Many also called for stronger emphasis on ”fundamentals,” ”critical thinking,” and ethical considerations:
+Despite overall confidence in their curriculum, in responding to the first institutional recommendations response question, over half of students wanted to be taught how to use GenAI.
+Many also called for stronger emphasis on fundamentals, critical thinking, and ethical considerations:
 “I think focusing on the idea of working alongside GenAI on how to take advantage of it to produce better, faster results, but acknowledge its limitations […] and discussing ethical and other ramifications of using AI in certain situations.
@@
 “Show the downsides and flaws of things like gpt-4, discuss environmental impacts, and don’t use it. GenAI is built on stolen data sets and harms artists. Stop endorsing its use. — Undergraduate Other
-The findings reveal a misalignment between perceived career impact and readiness.
 As GenAI begins to enter technical curricula, students in other disciplines risk being left behind.
-Addressing this gap will require expanding GenAI education across fields, embedding GenAI literacy into non-STEM programs, and equipping faculty to lead informed, inclusive discussions.
-These efforts are essential to preparing a more equitable and capable generation for a GenAI-driven future.
-Concern: learning at what cost?
-Student concerns about GenAI range from ethics and job security to institutional and faculty preparedness.
+This gap underscores a fundamental pedagogical challenge: while students recognize GenAI’s relevance, they often do not see it meaningfully integrated into their learning environments, leaving students to independently navigate tools that carry both potential and risk.
+This lack of structure points to an urgent need for faculty guidance and training, intentional curriculum design, and interdisciplinary GenAI literacy across academic programs.
+These efforts are essential to prepare a more equitable and capable generation for a GenAI-driven future.
+Concern: use at what cost?
+Student concerns about GenAI range from ethical implications and job security to institutional and faculty preparedness.
 While excitement is widespread, it is tempered by deeper anxieties, reflecting the complexity of the GenAI Generation’s experience.
@@
 I think GenAI can help me do what I like, which would benefit people and society.“ — Graduate Computer Science
-“im excited for the changes and technological advances it will bring.“ — Graduate Computer Science
+“I’m excited for the changes and technological advances it will bring.“ — Graduate Computer Science
 In contrast, positive outlook drops to 64% among undergraduate computer science students, reinforcing how exposure shapes perception.
@@
 “I think it is important to emphasize both positives and negatives of GenAI whenever mentioned. ” — Undergraduate Computer Science
-Free responses reflect strong concern, especially in the “additional thoughts” section.
-Over 50% conveyed negative emotions, with recurring themes of ethics, job replacement, and uncertainty—often expressed in long, passionate comments.
+Free responses reflect strong concern, especially in the the open ended qustion asking for additional thoughts.
+Over 50% of undergraduates conveyed negative emotions, with recurring themes of ethics, job replacement, and uncertainty—often expressed in long, passionate comments.
 “I think it will actively damage the quality of work and data produced[…]” — Undergraduate Engineer
@@
 “Goodbye middle class, we doomed.” — Undergraduate Computer Science
-“i’m scared it’s going to make my career not what i want it anymore.” — Undergraduate Other
+A compelling insight lies in who chose to share their perspectives in the optional free-response section.
+Among undergraduates, participation was highest among engineering and “other” majors, with response rates of 50% and 44%, respectively-compared to just 20% among computer science undergraduates.
+This could suggest that students who may not encounter GenAI in their formal coursework still feel strongly about its impact and relevance.
 Student responses reveal strong concerns about the ethical, economic, and societal impacts of GenAI, alongside a desire for greater guidance and critical engagement.
-Institutions must actively support both students and faculty as they navigate this rapidly evolving landscape.
+Students are not only grappling with how to use GenAI but also with what kind of future it is shaping.
 5 Conclusion
-The GenAI Generation, defined as students in their formative educational years during the rise of GenAI, face a rapidly evolving landscape shaped by both opportunity and uncertainty.
-While many are excited about GenAI’s potential, they also express concerns around preparedness, ethics, and long-term implications.
-Findings show that students with greater exposure to GenAI, particularly graduate computer science students, report higher confidence and optimism about their future careers.
-However, such structured exposure is not consistent across disciplines, revealing an institutional gap that must be addressed to ensure equitable support for all students navigating this technological shift.
-The GenAI Generation is distinct in its developmental context as students are not only adapting to new technologies but doing so without established norms or clear guidelines.
-They are the early adopters of tools like ChatGPT, learning to balance traditional skills with new competencies required to interact with GenAI effectively.
-Their adaptability is frequently tested while grappling with ethical questions, workforce changes, and broader societal implications.
-Students want more than just exposure: they seek guidance on using GenAI tools responsibly, critical thinking around AI-generated outputs, and a robust understanding of ethical considerations.
-The GenAI Generation is asking for a holistic approach that not only enhances technical skills but also nurtures critical engagement and ethical reasoning.
-It is critical to closely monitor and derive clarity on the potential impacts of these tools, developing a deeper understanding of cognition and ethical implications.
-The GenAI Generation is at a pivotal moment.
-As the first cohort to learn and grow in a GenAI-centric world, their experiences will set the tone for future generations.
-The data highlights the need for GenAI literacy efforts that span disciplines and prepare both students and faculty to engage with these tools in meaningful ways.
-The voices of these students are clear: they want to be part of a future where GenAI is a tool for growth, not a source of uncertainty, and where their education equips them to use GenAI as a force for positive change.
-Acknowledgements
-We would like to thank the students who took the time to respond to our survey, Dr.
-Joseph Ludlum and Dr.
-Loraine Phillips from the Office of Academic Effectiveness at Georgia Tech for their support in distributing the Fall 2023 cohort survey, and Dr.
-Thomas Ploetz for the continued encouragement throughout the process.
+This study contributes to the growing discourse on generative AI in education by examining how students are navigating both opportunity and uncertainty without consistent institutional guidance.
+It introduces the term GenAI Generation to describe those in their formative educational years during the rise of GenAI, students whose academic development is unfolding alongside the rapid adoption of these tools.
+Through analysis of over 250 survey responses, the findings reveal not only patterns of perceived readiness and ethical concern, but also a broader psychosocial landscape and pedagogical gap.
+Students are engaging with GenAI amid unclear norms, forming academic identities and career expectations as institutions lag behind.
+These insights offer timely direction for educators and policymakers seeking to design inclusive, developmentally aligned, and ethically grounded approaches to GenAI integration.
+The GenAI Generation faces a rapidly evolving landscape shaped by both excitement about GenAI’s potential and concern about preparedness, ethics, and long-term implications.
+This study reveals a growing divide in student confidence and readiness, shaped more by exposure and self-directed integration of GenAI than by formal curriculum.
+However, structured incorporation is inconsistent across disciplines, highlighting a broader institutional gap that must be addressed to ensure equitable support for all students navigating this technological shift.
+The divide is no longer simply between STEM and non-STEM fields, but between those with and without access to meaningful engagement with emerging technologies.
+The GenAI Generation is distinct in its developmental context.
+These students are not only adapting to powerful new tools, but doing so without established norms or guidance.
+As early adopters, they must balance traditional skills with new competencies while grappling with ethical dilemmas, shifting workforce expectations, and unknown cognitive impacts.
+Their adaptability is being tested in real time.
+Students want more than access, they are asking for structure.
+They seek guidance on using GenAI responsibly, developing critical thinking around AI-generated outputs, and building a strong foundation in ethical reasoning.
+Their feedback calls for a holistic approach that prepares them not only to use GenAI, but to question and shape it.
+As the first cohort to learn and grow in a GenAI-centric world, the experiences of the GenAI Generation will set the tone for future educational trajectories.
+Their voices point to an urgent need for GenAI literacy that spans disciplines and prepares both students and faculty to engage with these tools thoughtfully and equitably.
+They envision a future where GenAI is not a source of confusion or inequity, but a tool for growth—integrated into education in ways that are ethical, empowering, and human-centered.
+6 List of abbreviations
+Abbreviation Full Form
+GenAI Generative Artificial Intelligence
+AI Artificial Intelligence
+IRB Institutional Review Board
+OMSCS Online Master of Science in Computer Science
+MSCS Master of Science in Computer Science
+OAE [Institution Name Redacted for Review] Office of Academic Effectiveness
+STEM Science, Technology, Engineering, and Mathematics
+CS Computer Science (used in tables/figures for grouping)
+7 Future Considerations
+To better understand the evolving impact of GenAI on student development and educational outcomes, future research should consider longitudinal studies that track student perceptions, readiness, and concerns over time.
+As GenAI tools rapidly advance, students’ experiences and attitudes will likely shift in response to both technological innovation and institutional adaptation.
+Capturing these changes across semesters or academic years would provide valuable insight into how exposure, policy, and pedagogy shape preparedness and perception in the long term.
+In addition to tracking perception, researchers should examine the cognitive consequences of increased reliance on GenAI tools, such as ChatGPT, for learning and problem-solving.
+Cognitive offloading may free students to focus on higher-order thinking, but it also raises questions about memory retention, deep learning, and critical reasoning.
+Understanding how sustained use of GenAI affects cognitive engagement, information retention, and academic identity will be critical to developing pedagogical strategies that support—not replace—core learning processes.
+There is also a need for targeted dissemination and outreach to disciplines beyond computer science.
+This study found that students in non-computer science fields, particularly those in the ”Other” category, demonstrated strong emotional responses and a desire for inclusion in GenAI discourse, despite limited curricular exposure.
+Expanding GenAI literacy and integration into these areas can help bridge the readiness gap and empower students across all domains to engage critically and confidently with this transformative technology.
+8 Declarations
+8.1 Funding
+This study was conducted without external funding. No funding agency was involved in the design, data collection, analysis, or publication of this research.
+8.2 Funding Statement
+8.3 Consent to Publish declaration
+Not applicable
+8.4 Consent to Participate
+This research study was Institutional Review Board (IRB) approved. Respondents provided informed consent.
+8.5 Ethics declaration
+This research study was Institutional Review Board (IRB) approved.
+The study collected students’ perceptions of their future career preparedness in light of the potential impact of generative artificial intelligence (GenAI) through a survey.
+The survey only collected responses for students located within the United States, and 18 years of age or older as per IRB instruction.
+Respondents provided informed consent.
+Survey questions were designed to protect participant identity.
+Participation was completely voluntary, and had the ability to opt out at any time.
+No benefit or compensation was provided.
+We complied with any applicable laws and regulations regarding confidentiality.
+The IRB and Office of Human Research Protections reserved the right to look at study records at any time.
+Contacts to the author and Office of Research Integrity Assurance were provided.

### v3

@@
 Generative Artificial Intelligence (Gen AI) is revolutionizing education and workforce development, profoundly shaping how students learn, engage, and prepare for their future.
-Outpacing the development of uniform policies and structures, GenAI has heralded a unique era and given rise to the GenAI Generation.
-We define the GenAI Generation as a cohort of students whose education has been increasingly shaped by the opportunities and challenges GenAI presents during its widespread adoption within society.
-This study examines students’ perceptions of GenAI through a concise survey with optional open-ended questions, focusing on their awareness, preparedness, and concerns.
-Notably, readiness appears increasingly tied to exposure to GenAI through one’s coursework.
-Students with greater curricular exposure to GenAI tend to feel more prepared, while those without it more often express vulnerability and uncertainty, highlighting a new and growing divide in readiness that goes beyond traditional disciplinary boundaries.
-Evaluation of more than 250 responses, with over 40% providing detailed qualitative feedback, reveals a core dual sentiment: while most students express enthusiasm for GenAI, an even greater proportion voice a spectrum of concerns about ethics, job displacement, and the adequacy of educational structures given the highly transformative technology.
+Outpacing the development of uniform policies and structures, Gen AI has heralded a unique era and given rise to the Gen AI Generation, a cohort of individuals whose development has been increasingly shaped by the opportunities and challenges Gen AI presents during its widespread adoption within society.
+This study examines higher education students’ perceptions of Gen AI through a concise survey with optional open-ended questions, focusing on their awareness, preparedness, and concerns.
+”Readiness” appears increasingly tied to exposure to Gen AI through one’s coursework.
+Students with greater curricular exposure to Gen AI tend to feel more prepared, while those without more often express vulnerability and uncertainty, highlighting a new and growing divide that goes beyond traditional disciplinary boundaries.
+Evaluation of more than 250 responses, with over 40 percent providing detailed qualitative feedback, reveals a core dual sentiment: while most students express enthusiasm for Gen AI, an even greater proportion voice a spectrum of concerns about ethics, job displacement, and the adequacy of educational structures given the highly transformative technology.
 These findings offer critical insights into how students view the potential and pitfalls of Gen AI for future career impacts.
-The challenge ahead involves implementing associated recommendations for educational institutions, moving beyond the baseline of access toward more informed guidance on the use of these tools, while preserving critical thinking, ethical reasoning, and adaptive learning.
-Keywords: Generative Artificial Intelligence; GenAI Generation; Student Perceptions; Higher Education; Career Readiness; Cognitive Development; Adaptability; Problem-Solving; Critical Thinking; Ethics in AI
-1 Introduction
-The significant advancement of Artificial Intelligence (AI), particularly Generative AI (GenAI), is reshaping the educational landscape.
-GenAI is a type of AI trained on large datasets to generate new content such as text, images, audio, or video in response to user prompts, setting it apart from predictive models by its ability to create [22, 28, 20, 21].
-As a tool capable of autonomously producing human-like content, GenAI offers both opportunities and uncertainties for education.
-It can enhance creativity, streamline workflows, and support personalized learning, yet its broader impact remains unclear.
-As GenAI continues to evolve exponentially and integrate within academic contexts, there is a pressing need to understand how it is shaping the developmental experiences of students in their formative academic years.
-Unlike previous waves of technology’s impact on education, GenAI’s pace and scale introduce a layer of uncertainty around its long-term effects on cognition, critical thinking, and engagement.
-By honing in on student perspectives, this research provides early insight into how GenAI is influencing not only perceived career readiness but also the cognitive processes underpinning learning and decision-making.
-This study presents the development, dissemination, and evaluation of a survey examining student perspectives on GenAI through three core lenses: awareness, preparedness, and concern.
-By analyzing responses, it provides actionable insights into how the GenAI Generation—students most directly impacted by the rise of GenAI—understand its role in shaping both their academic experiences and future trajectories.
-The GenAI Generation
+I Introduction
 The Gen AI Revolution represents a paradigm shift in technology adoption, enhancing efficiency, creativity, and performance while introducing challenges such as ethical concerns and the need for responsible stewardship [12, 5, 17, 4].
-As it relates to education, this revolution not only emphasizes a transformation in learning but also highlights the challenges students and teachers face during a phase of unclear uniform guidance [4, 5].
 This study introduces the Gen AI Generation.
-The GenAI Generation is defined as the cohort of individuals actively undergoing developmental and educational growth/attainment during the rise of Generative AI, starting with the release of ChatGPT in November 2022.
-The GenAI Generation encompasses the experiences of students from high school through higher education, during the initial, transformative years of the GenAI Revolution, when the integration of AI tools is still experimental, unregulated, and highly influential.
+The Gen AI Generation is defined as the cohort of individuals actively undergoing developmental and educational growth/attainment during the experimental, unregulated rise of Gen AI.
 Unlike earlier generations who did not have these tools and later generations who may inherit more structured systems, this group faces unique challenges and opportunities surrounding uncertainty.
@@
 Long-Term Impact Unknown: Using these tools during formative years without a clear understanding of potential long-term cognitive or social consequences.
-Figure 1: Illustrates the relationships between key characteristics, emphasizing the ongoing feedback loop that drives development and adaptation amidst uncertainty.
-Just as Google transformed how students accessed and retained information, leading to phenomena dubbed the “Google Effect” [26], and the rise of digital technologies sparked concerns about “digital dementia” [27], the GenAI Revolution represents another critical shift in the educational landscape.
-These past transformations demonstrate how new technologies can reshape cognitive processes, social behaviors, and academic growth, sometimes with unintended, long-term consequences [25, 6, 18].
-The GenAI Generation, defined by its unique characteristics, is similarly at the forefront of this era’s challenges and opportunities.
-Evaluating this cohort is essential to understand how these tools influence learning and development during critical formative years, ensuring that we can build a foundation of guidance, policies, and research to support them and future generations effectively.
+Just as Google transformed how students accessed and retained information, dubbed the “Google Effect” [22], and the rise of digital technologies sparked concerns about “digital dementia” [23], the Gen AI Revolution represents another critical shift.
+Evaluating the Gen AI Generation is essential to understand how these tools influence development, ensuring we can build a foundation of guidance and informed policies.
 Awareness of Gen AI’s Capabilities and Applications
-For the purpose of this study, students’ awareness of GenAI means their familiarity with tools such as ChatGPT and their understanding of how these tools can be leveraged in educational and professional contexts.
-Research has revealed varying levels of awareness globally.
-One study found widespread awareness among students across 76 countries, with significant usage for academic tasks such as information retrieval and paraphrasing, though concerns about academic integrity and creativity persist [29].
-Another study highlighted differences in engagement with GenAI based on demographic factors, underscoring the need for targeted educational support [10].
-Research also shows that while GenAI is valued for idea generation and brainstorming, there are reservations regarding its use in completing assignments, emphasizing the need for ethical policies [3].
-The transformative potential of GenAI in education has been widely discussed, particularly in its applications for personalized learning and adaptive tutoring systems, which support higher-order cognitive tasks such as problem-solving and critical analysis [8].
-However, challenges like data privacy and instructor resistance remain significant barriers to effective integration.
-Other studies have explored how personalized learning and language support tools can aid students while also warning of potential over-reliance and its impact on critical thinking [17].
-Current research on student awareness of GenAI’s impacts reveals perspectives shaped by demographic and geographic factors.
-These findings highlight the importance of adopting diverse approaches to educate students about GenAI while providing guidance on data privacy and ethical considerations.
+For this study, students’ awareness of GenAI refers to their familiarity with Gen AI tools and their understanding of how such tools can support academic and professional work.
+Prior research shows high global awareness and frequent use for tasks such as information retrieval and paraphrasing, though concerns about academic integrity and creativity remain [24].
+Engagement with GenAI also varies across demographic groups, indicating a need for targeted educational support [11].
+While students value GenAI for idea generation and brainstorming, many remain hesitant to use it for completing assignments, underscoring the need for clear ethical guidance [3].
+Work on GenAI’s educational potential highlights its promise for personalized learning and adaptive tutoring that foster higher-order cognitive skills [7].
+Additional studies note that while personalized support tools can aid learning, over-reliance may hinder critical thinking, reinforcing the need for balanced instructional approaches [18].
+Current research shows student awareness is shaped by geo-demographic factors, suggesting that diverse and ethically grounded strategies are needed to prepare students for GenAI’s role in education.
 Preparedness for an Gen AI-Driven Workforce
-For the purpose of this study, student preparedness refers to their confidence in navigating a dynamic, AI-driven job market and their perception of how well their education equips them for this future.
-Research highlights gaps in AI-specific training within higher education curricula, which leave students feeling under-prepared for the evolving workforce [1, 7].
-Ethical and accuracy concerns are also noted as barriers to effective integration, further underscoring the importance of incorporating AI literacy and practical applications into academic programs to enhance student readiness [7].
-Additionally, project-based and interdisciplinary learning approaches are recommended to prepare students to critically engage with AI tools and align their academic training with workforce demands [9].
-Other studies propose professional development programs for educators to ensure that GenAI integration aligns with institutional goals and emphasizes critical digital literacy, enabling both students and faculty to navigate the complexities of GenAI in professional environments [23].
-Personalized learning supported by GenAI is also identified as a potential way to improve academic and employability outcomes, though concerns about ethical use and critical thinking remain [12].
-Students’ preparedness highlights the interdisciplinary importance of addressing GenAI in the classroom, ensuring clarity and intentionality in these discussions.
-Although this research does not examine faculty perspectives, it is critical to recognize that educators also require support and guidance to effectively integrate GenAI discussions and tools into their curricula.
+For this study, student preparedness refers to their confidence in navigating an AI-driven job market and their sense of how well their education equips them for that future.
+Research identifies persistent gaps in AI-focused training within higher education, leaving many students feeling under-prepared, especially amid ethical and accuracy concerns that hinder wider integration [1, 6].
+Scholars recommend embedding AI literacy, practical applications, and interdisciplinary, project-based learning to better align academic preparation with workforce demands [8].
+Faculty readiness is also critical: professional development programs can help ensure GenAI adoption supports institutional goals and strengthens critical digital literacy for both educators and students [20].
+Personalized learning supported by GenAI may further enhance academic and employability outcomes, though concerns about ethics and impacts on critical thinking remain [13].
+Overall, student preparedness underscores the need for intentional, interdisciplinary engagement with GenAI in the classroom, alongside adequate support for educators integrating these tools.
 Concerns About Gen AI’s Potential Implications
-For the purpose of this study, students’ concerns refer to their anxiety or worry about GenAI’s potential impact.
-Current research highlights ethical issues such as data privacy and bias, alongside practical challenges like overreliance and the risk of misinformation.
-Studies have noted strong reservations about using GenAI for completing assignments without disclosure as well as concerns about the uncritical acceptance of AI outputs [3, 2].
-Additionally, research underscores how the misuse of GenAI could undermine academic integrity and originality, emphasizing the need for critical evaluation skills and clear institutional guidelines for responsible use [29, 2].
-Balancing the benefits of automation with safeguards against biases and ethical risks remains a pressing concern.
-Some researchers critique traditional assessment models, advocating for redesigned evaluations that prioritize originality and critical engagement [19].
-Others warn that over-reliance on GenAI could diminish critical thinking and intellectual engagement, urging for its balanced integration into educational practices [17].
-While GenAI tools have the potential to enhance advanced cognitive processes, concerns persist regarding their impact on foundational learning skills such as critical thinking and engagement [3, 29, 2, 19, 17].
-This contrast highlights the urgent need for clear guidance on student use of GenAI across all educational levels.
-Strategies must be developed to ensure that these tools are employed ethically while fostering constructive critical engagement and safeguarding cognitive development.
+For this study, students’ concerns refer to their anxiety or worry about GenAI’s potential impact.
+Research highlights ethical risks, such as data privacy, bias, and misinformation, alongside worries about overreliance and uncritical acceptance of AI outputs [3, 2].
+Students express particular reservations using GenAI for assignments without disclosure, and studies warn that misuse may undermine academic integrity and originality, underscoring the need for critical evaluation skills and clear institutional guidelines [24, 2] [19].
+Related work warns that over-reliance on GenAI could weaken critical thinking and foundational learning skills, calling for balanced and intentional integration into educational practices [18].
+While GenAI can support advanced cognitive processes, persistent concerns about ethics, bias, and cognitive impacts highlight the need for coherent guidance that promotes responsible use while safeguarding student development [3, 24, 2, 19, 18].
 Contributions of This Study
-This study offers timely insights into how students in their formative educational years are experiencing the rapid rise of Generative AI.
-It introduces the term “GenAI Generation” to describe students navigating both academic development and technological transformation without consistent institutional guidance or established norms.
-Drawing from a combination of structured survey data and rich qualitative responses, the study reveals meaningful patterns in students’ awareness, preparedness, and concern regarding GenAI.
-Across majors, students expressed a desire for more intentional support, not just access to GenAI tools, but structured opportunities to learn how to use them responsibly.
-While formal curriculum integration remains inconsistent, students who encounter GenAI through coursework or self-directed exploration are more likely to articulate confidence and optimism about its role in their future.
-From enthusiasm to anxiety, the findings illuminate the psychosocial and pedagogical dimensions of GenAI’s presence in higher education.
-Students are not only engaging with a new technology, they are forming perceptions of their own readiness, trust in institutions, and ethical responsibility in real time.
-These perspectives carry clear implications for curriculum design, faculty preparation, and equitable access to digital literacy.
-This work contributes to an emerging understanding of what students need to thrive in a GenAI-influenced world: not only technical fluency, but critical engagement, ethical grounding, and institutional clarity.
-The study underscores the importance of proactive educational responses that extend beyond STEM fields to support students across all disciplines in this pivotal moment.
-2 Methods
-The survey targeted students within the GenAI Generation, focusing on GenAI’s perceived impact on their education and careers.
-Awareness, preparedness, and concern guided the development of eight main questions, three demographic items, and two open-ended prompts.
-This study received Institutional Review Board (IRB) approval from the [Institution Name Redacted for Review], and all participants provided informed consent prior to beginning the survey.
-Designed for completion in under five minutes [13], the survey began with an IRB-required disclosure and a screening question, followed by items progressing from general to sensitive topics [24].
-Questions were refined through lab feedback, aligned with key concepts, and included internal validation to ensure consistency and quality.
-A four-point, forced-choice Likert-type scale was used to assess attitudes, significance, intensity, and satisfaction.
-While omitting a neutral option is debated [14], this was an intentional design choice to encourage engagement and reduce bias [15].
-The survey was administered across two [Institution Name Redacted for Review] student cohorts, with full question text available in the supplemental materials.
-The first cohort, Fall 2023 Computer Science Graduate students, survey was distributed to Online Computer Science Master’s (OMSCS) students, approximately 12,000 members [CITE REDACTED FOR REVIEW], and campus Computer Science Master’s (MSCS), approximately 14,500 members [CITE REDACTED FOR REVIEW], via listservs.
-Invitations were sent on September 28 and 29, 2023, with responses collected through November 1.
-The second cohort, Fall 2023 Undergraduate students, survey was conducted with support from the [Institution Name Redacted for Review] Office of Academic Effectiveness (OAE), which used stratified random sampling to select 5,655 undergraduates representative of the broader student population.
-Sampling quotas ensured balance by gender, race, citizenship, and class standing, with a 2% margin of error.
-The demographic breakdown of the sample can be found in the supplemental materials.
-The survey was open from November 2–23, 2023, with email reminders sent on November 2, 15, and 22.
-Analysis Approach
-Descriptive statistics and frequency distributions were used to compare response patterns across student subgroups, while qualitative evaluation of open-ended responses enabled a systematic exploration of emerging themes and perceptions.
-Data Processing: To ensure data integrity, only completed responses containing degree type information were included in the analysis.
-Where the first cohort of graduate students comprised soley of computer science majors, the undergraduate cohort had many degrees.
-Rather than broadly distinguishing between STEM and non-STEM, degrees were categorized into Computer Science, Engineering, and Other, based on disciplinary focus and methodological orientation.
-Table 1 provides a Major-Category mapping for reference.
-This grouping structure reflects both practical sample sizes (Other, 28 Computer Science, and 60 Engineering) and a meaningful distinction in how students may engage with GenAI in academic contexts.
-This also enabled a comparison between computer science graduate and undergraduate perspectives.
-Grouping a diverse range of degrees—such as Biology, Business Administration, Literature, and Public Policy—under the single category “Other” introduces a potential limitation, as it may obscure important disciplinary differences in students’ exposure to or perspectives on GenAI.
-While specific exposure to AI tools and curricula was not measured directly, these categories serve as a reasonable proxy for varying degrees of disciplinary integration with GenAI-related content, allowing exploratory comparisons across educational domains.
-Table 1: Undergraduate Sample and Response Rates by Majors and Category
-Major* Category Sample Freq. Response Freq. Sample % Response %
-Applied Physics Other 12 2 0.2 1.7
-Biology Other 162 3 2.9 2.6
-Business Administration Other 440 3 7.8 2.6
-Computational Media Other 98 1 1.7 0.9
-Earth & Atmospheric Sciences Other 13 1 0.2 0.9
-Econ & Int’l Affairs Other 15 1 0.3 0.9
-Economics Other 49 2 0.9 1.7
-History, Technology, & Society Other 18 2 0.3 1.7
-Industrial Design Other 88 2 1.6 1.7
-Int’l Affairs & Mod Lang Other 18 2 0.3 1.7
-Lit., Media, & Communication Other 59 2 1.0 1.7
-Mathematics Other 77 3 1.4 2.6
-Neuroscience Other 146 3 2.6 2.6
-Physics Other 68 5 1.2 4.3
-Psychology Other 66 5 1.2 4.3
-Public Policy Other 37 1 0.7 0.9
-Aerospace Engineering Engineer 349 8 6.2 7.0
-Biomedical Engineering Engineer 399 6 7.1 5.2
-Chemical and Biomolecular Eng Engineer 187 5 3.3 4.3
-Computer Engineering Engineer 187 2 3.3 1.7
-Electrical Engineering Engineer 187 3 3.3 2.6
-Environmental Engineering Engineer 58 3 1.0 2.6
-Industrial Engineering Engineer 344 5 6.1 4.3
-Materials Science & Engr Engineer 67 3 1.2 2.6
-Mechanical Engineering Engineer 582 12 10.3 10.4
-Nuclear & Radiological Engr Engineer 32 2 0.6 1.7
-Computer Science Computer Science 1419 28 25.1 24.3
-Applied Lang/Intercultural St No Response 4 0 0.1 0.0
-Architecture No Response 76 0 1.3 0.0
-Atmospheric & Oceanic Sciences No Response 3 0 0.1 0.0
-Biochemistry No Response 88 0 1.6 0.0
-Building Construction No Response 13 0 0.2 0.0
-Chemistry No Response 40 0 0.7 0.0
-Environmental Science No Response 10 0 0.2 0.0
-Global Econ/Mod Lang No Response 5 0 0.1 0.0
-International Affairs No Response 22 0 0.4 0.0
-Music Technology No Response 23 0 0.4 0.0
-Solid Earth/Planetary Sciences No Response 2 0 0.0 0.0
-Special/Non-Degree No Response 64 0 1.1 0.0
-Civil Engineering No Response 128 0 2.3 0.0
-*For reference; sample was not balanced on this attribute. Sample breakdowns by additional attributes can be found within the supplemental material.
-For the undergraduate cohort, two non-undergraduate entries, likely from dual-enrolled students, were excluded to maintain a focus on undergraduate perspectives.
-The remaining responses totals 102 for the graduate cohort and 115 for the undergraduate cohort.
-Table 2 shows this process
-Table 2: Table showing data cleaning process
-Survey Population Total Started Total Incomplete Total Dropped Reason for Drop Remaining Completed Total
-Computer Science Graduates 2023 26,000** 122 19 1 Response was “prefer not to answer” for degree 102
-Undergraduates 2023 5,655 143 26 2 Did not have “undergraduate” for degree type 115
-** Approximate total population estimate for Computer Science Graduate 2023 cohort.
-Qualitative Free-Response Analysis: An inductive coding approach was used to analyze free-response data.
-Preprocessing included tokenization, lowercasing, and the removal of punctuation and stop words.
-N-grams (unigrams to quadgrams) were extracted, with low-frequency patterns excluded.
-Grammatical parsing identified noun and verb phrases, which, alongside n-grams, helped surface recurring themes and patterns.
-3 Results
+This study provides insight into how higher education students are experiencing the rapid rise of Gen AI, introducing the term “Gen AI Generation”.
+Drawing on structured survey data and qualitative responses, the findings reveal patterns in students’ awareness, preparedness, and concern, showing a clear desire for intentional support.
+II Methods
+The survey included eight main questions, three demographic items, and two open-ended prompts.
+The study received Georgia Tech (GT) IRB exemption, and all participants provided informed consent.
+Designed for completion in under five minutes [14], the survey opened with required disclosures and a screening item, progressing from general to more sensitive topics [21].
+Items were refined through lab feedback and internal validation.
+Attitudes were measured using a four-point forced-choice Likert-type scale, intentionally omitting a neutral option to encourage engagement and reduce satisficing [15, 16].
+Survey administration and full question wording are provided in the supplemental materials.
+Two GT cohorts were sampled.
+The Fall 2023 Computer Science Graduate cohort included online ($\approx$12,000 students[10]) and on-campus students ($\approx$14,500[9]), recruited via listservs from Sept.
+28–Nov.
+1, 2023.
+The Fall 2023 Undergraduate cohort was recruited through the Office of Academic Effectiveness using stratified random sampling of 5,655 undergraduates balanced by gender, race, citizenship, and class standing (2% margin of error).
+The survey ran Nov.
+2–23, 2023 with reminder emails on Nov.
+2, 15, and 22.
+Demographic breakdowns appear in the supplemental materials.
+Analysis included descriptive statistics and frequency distributions to compare subgroup patterns, supplemented by qualitative analysis of open-ended responses.
+For qualitative data, preprocessing involved tokenization, lowercasing, removal of punctuation and stop words, and extraction of unigrams through quadgrams with low-frequency terms excluded.
+Grammatical parsing of noun and verb phrases supported inductive coding to identify recurring themes.
+III Results
 The undergraduate cohort included 115 respondents across three categories: 28 in Computer Science, 60 in Engineering, and 27 in Other disciplines.
 The graduate cohort response count was 102 and comprised solely of computer science majors.
-The full breakdown between sample to response counts can be seen in Table 1.
+The full breakdown between sample to response counts can be seen in the supplemental materials.
 Figure 1: Response counts by degree and level
+Figure 2: Responses on Gen AI awareness.
 Awareness: A small percentage of respondents in all groups believe Gen AI is a short-lived trend, seen in Figure 2, with undergraduate computer science students slightly more likely to agree at 21%.
-Most respondents view generative AI as a significant advancement, with nearly unanimous agreement among the graduate cohort at 100%, followed by undergraduate computer science and engineers at 96% and 93%, respectively.
-When asked about the limitations of generative AI’s potential applications, agreement levels vary, with engineers and graduate respondents showing slightly higher agreement at 34% and 37%.
-Generative AI experience is highest among Graduate respondents (93%) and lowest among Engineers (78%).
-Figure 3: Responses on GenAI awareness.
+Most respondents view Gen AI as a significant advancement, with nearly unanimous agreement among the graduate cohort at 100%, followed by undergraduate computer science and engineers at 96% and 93%, respectively.
+When asked about the limitations of Gen AI’s potential applications, agreement levels vary, with engineers and graduate respondents showing slightly higher agreement at 34% and 37%.
+Gen AI experience is highest among Graduate respondents (93%) and lowest among Engineers (78%).
 Preparedness: For the percentage of respondents that believe job skills will change due to Gen AI, Figure 3 shows undergraduate computer science students have the highest percentage at 82%, followed by computer science graduates at 77%.
@@
 Concern: The majority view Gen AI’s career impact positively, seen in Figure 6, with computer science graduates highest at 81% and undergraduate Engineers, Computer Science, and Other lower with 70%, 64%, and 63% respectively.
-Excitement about GenAI benefits was high across all groups, peaking among Engineering at 77% and Test at 80%.
+Excitement about Gen AI benefits was high across all groups, peaking among Engineering at 77% and Computer Science Grads at 80%.
 Other respondents notably lower at 67%.
@@
 Free Response: In addition to the structured survey responses, a notable portion of participants contributed to the optional open-ended questions.
-These qualitative insights deepen understanding of the perspectives behind the patterns and are summarized below.
+These qualitative insights deepen understanding of the perspectives behind the patterns.
 40–50% of undergraduates and graduate students responded to the institutional recommendation prompt, and 23–25% provided input on the additional thoughts prompt.
-Participation rates for the free response can be seen in Table 3, showing non-Computer Science students having the highest participation.
-The Fall undergraduate free responses were analyzed as a whole.
-Table 3: Optional Free Response Participation by Question & Degree Category
-Category Institutional Recommendation Total Additional Thoughts Total Survey Response Total
-Undergraduate Other 17 11 38
-Undergraduate Engineer 24 11 49
-Undergraduate Computer Science 6 5 29
-Graduate Computer Science 52 27 102
 Figure 7: Topics raised in free response for institutional recommendations
-The most common recommendation, mentioned by over 50% for both cohorts, was teaching students how to use GenAI tools (Figure 8.
+The most common recommendation, mentioned by over 50% for both cohorts, was teaching students how to use Gen AI tools (Figure 7).
 About 33% of respondents also suggested integrating Gen AI into existing curricula or adding new courses.
 Ethics and career implications of Gen AI appeared more frequently in undergraduate responses, with ethics mentioned 15% more and career impact 5% more than in graduate responses.
-There were fewer who responded to the optional ”additional thoughts” question, those who did showed clear emotional tone(Figure 9).
-Over half of undergraduates expressed negative emotions, 29% more than the graduate cohort, while graduates were 20% more likely to express positivity.
+There were fewer who responded to the optional additional thoughts question, those who did showed clear emotional tone(Figure 8).
+Over half of undergraduate free responses expressed negative emotions, 29% more than the graduate cohort, while graduates were 20% more likely to specifically express positivity.
 Additional thoughts covered a broader range of themes(Figure 9).
@@
 While the first prompt focused on Gen AI’s utility, the second invited any remaining reflections from participants.
-4 Discussion
+IV Discussion
 The results have presented unique and shared perspectives across undergraduate programs, revealing areas of improvement.
-Thanks to the participants in this study and their contributions to the free response, there is a story that has begun to emerge surrounding the needs of the GenAI Generation as they learn and grow into their future careers.
-Although surveys were distributed uniformly to ensure equal access across cohorts, the potential for selection bias and low response frequencies remains a limitation.
-Response counts were modest compared to sample size, and some subgroups—particularly within non-CS or less tech-aligned majors—were underrepresented.
-As a technology-focused institution, [Institution Name Redacted for Review]’s student population may also have greater exposure to emerging technologies, including GenAI, than students at other institutions.
+Although surveys were distributed uniformly, the potential for selection bias and low response frequencies remains a limitation.
+Response counts were modest compared to sample size, and some subgroups, particularly within non-STEM majors, were underrepresented.
+As a technology-focused institution, Georgia Tech’s student population may also have greater exposure to emerging technologies, including Gen AI, than students at other institutions.
 Even with these constraints, the findings offer important early insight into how Gen AI is shaping student experiences, perceptions, and expectations during a critical period of educational and technological transformation.
 Awareness: the uncertainty in early adopters
-While overall awareness of GenAI was high across all groups, levels of understanding and confidence varied.
-Computer science undergraduates were the most likely to consider GenAI a short-lived trend (21%), whereas engineering and “Other” majors reported lower agreement (10–11%).
+While overall awareness was uniformly high, levels of understanding and confidence varied.
+Computer science undergraduates were the most likely to consider Gen AI a trend (21%), whereas engineering and “Other” majors reported lower agreement (10–11%).
 Despite this skepticism, nearly all respondents across disciplines viewed Gen AI as a meaningful advancement, with agreement ranging from 89% to 100%.
@@
 I don’t feel like Gen AI has any negative impacts on my future.” — Undergraduate Engineer
-“I think it has limited application in anything with more advanced logic or understanding.” — Undergraduate Engineer
 “I think it has limited potential because Gen AI currently cannot distinct whether the information it gives is true or not.” — Undergraduate Engineer
 “It’s very uncertain and that’s what makes me slightly afraid. I also don’t understand it very well, which I should work on.” — Undergraduate Computer Science
-As seen in some of these responses, there is skepticism of GenAI achieving beyond what the current potential shows, lending to an experience of technology moving faster beyond students’ grasp.
+As seen in some of these responses, there is skepticism of Gen AI achieving beyond what the current potential shows, lending to a potential experience of technology moving faster beyond students’ grasp.
 Many students rely on self-directed learning, coursework, and personal experiences to form their perspectives on Gen AI.
-This gap in structured AI education leads to a divergence in awareness, confidence, and preparedness, with some students feeling optimistic and others expressing deep concern and uncertainty.
+This gap in structured AI education can lead to a divergence in awareness, confidence, and preparedness, with some students feeling optimistic and others expressing deep concern and uncertainty.
 In the optional free responses, 50% of the students expressed negative emotions, including frustration, fear, and worry, while 20% conveyed uncertainty, with overlapping sentiments highlighting the complex emotional landscape students navigate:
@@
 There is a clear pattern that begins to emerge: those who feel prepared in their degree are most likely to expect a positive career impact.
-Where computer science 77% of graduates and 82% undergraduates anticipate job skill changes, only 56% and 38% of undergaduate other and engineering students anticipate such changes, with many engineering students expecting a slower pace—highlighting a lack of urgency in their field.
+77% of computer science graduates and 82% computer science undergraduates anticipate job skill changes, only 56% and 38% of undergaduate other and engineering students anticipate such changes, with many engineering students expecting a slower pace, highlighting a lack of urgency in their field.
 These differences suggest the need for tailored strategies that align with each group’s career context.
-Despite these varied views, all cohorts reported relatively high confidence in their curricula—possibly reflecting trust in their degree or Institutions’ reputation.
+Despite these varied views, all cohorts reported relatively high confidence in their curricula, possibly reflecting trust in their degree or Institutions’ reputation.
 Still, engineering students stand out for their skepticism: they are the least likely to expect Gen AI-driven skill changes, report the lowest satisfaction with classroom discussions, and express the least confidence in faculty engagement.
@@
 The free response data adds depth to these findings.
-Despite overall confidence in their curriculum, in responding to the first institutional recommendations response question, over half of students wanted to be taught how to use GenAI.
+Despite overall confidence in their curriculum, over half of the free responses wanted to be taught how to use Gen AI.
 Many also called for stronger emphasis on fundamentals, critical thinking, and ethical considerations:
@@
 Until stronger evidence comes to light for its all-encompassing power, it should be ignored in higher education.” — Undergraduate Engineer
-“Show the downsides and flaws of things like gpt-4, discuss environmental impacts, and don’t use it. GenAI is built on stolen data sets and harms artists. Stop endorsing its use. — Undergraduate Other
-As GenAI begins to enter technical curricula, students in other disciplines risk being left behind.
-This gap underscores a fundamental pedagogical challenge: while students recognize GenAI’s relevance, they often do not see it meaningfully integrated into their learning environments, leaving students to independently navigate tools that carry both potential and risk.
+“Show the downsides and flaws of things like gpt-4, discuss environmental impacts, and don’t use it.
+Gen AI is built on stolen data sets and harms artists.
+Stop endorsing its use.
+— Undergraduate Other
+As Gen AI begins to enter technical curricula, students in other disciplines have an even greater risk being left behind.
+This gap underscores a fundamental pedagogical challenge: while students recognize Gen AI’s relevance, meaningful integration into their learning environments is falling behind, leaving students to independently navigate tools that carry both potential and risk.
 This lack of structure points to an urgent need for faculty guidance and training, intentional curriculum design, and interdisciplinary Gen AI literacy across academic programs.
@@
 Student concerns about Gen AI range from ethical implications and job security to institutional and faculty preparedness.
-While excitement is widespread, it is tempered by deeper anxieties, reflecting the complexity of the GenAI Generation’s experience.
-The Test 2023 Computer Science Graduate cohort reported the lowest concern (65%) and the highest levels of excitement (80%) and belief in GenAI’s positive career impact (81%).
+While excitement is widespread, it is tempered by anxieties, reflecting the complexity of the Gen AI Generation’s experience.
+The 2023 Computer Science Graduate cohort reported the lowest concern (65%) and the highest levels of excitement (80%) and belief in Gen AI’s positive career impact (81%).
 This likely reflects the integration of Gen AI into their curriculum and their confidence using such tools.
@@
 Even when students are unsure about Gen AI’s direct impact on their careers, many still express enthusiasm about its broader potential.
-Some specific positive takes included:
 “I think less people should have to work for as long and more leisure and volunteering time will be helpful for mental health and balance.” — Undergraduate Engineer
@@
 “I think it is important to emphasize both positives and negatives of Gen AI whenever mentioned. ” — Undergraduate Computer Science
-Free responses reflect strong concern, especially in the the open ended qustion asking for additional thoughts.
-Over 50% of undergraduates conveyed negative emotions, with recurring themes of ethics, job replacement, and uncertainty—often expressed in long, passionate comments.
+Free responses reflect strong concern, especially in the open ended question additional thoughts prompt.
+Over 50% of undergraduates conveyed negative emotions, with recurring themes of ethics, job replacement, and uncertainty, often expressed in long, passionate comments.
 “I think it will actively damage the quality of work and data produced[…]” — Undergraduate Engineer
@@
 Or maybe just politicians with bot accounts blindly supporting them and arguing with dissenters on social media.” — Undergraduate Engineer
-And some of the more succinct, poignant responses:
+And the more succinct, poignant response:
 “Goodbye middle class, we doomed.” — Undergraduate Computer Science
 A compelling insight lies in who chose to share their perspectives in the optional free-response section.
-Among undergraduates, participation was highest among engineering and “other” majors, with response rates of 50% and 44%, respectively-compared to just 20% among computer science undergraduates.
+Among undergraduates, participation was highest among engineering and “other” majors, with response rates of 50% and 44%, respectively-compared to just 20% among computer science undergraduate respondents.
 This could suggest that students who may not encounter Gen AI in their formal coursework still feel strongly about its impact and relevance.
 Student responses reveal strong concerns about the ethical, economic, and societal impacts of Gen AI, alongside a desire for greater guidance and critical engagement.
-Students are not only grappling with how to use GenAI but also with what kind of future it is shaping.
-5 Conclusion
-This study contributes to the growing discourse on generative AI in education by examining how students are navigating both opportunity and uncertainty without consistent institutional guidance.
+Students are not only grappling with how to use Gen AI, but also with what kind of future it is shaping.
+V Conclusion
+This study contributes to the growing discourse on generative artificial intelligence in education by examining how students are navigating both opportunity and uncertainty without consistent institutional guidance.
 It introduces the term Gen AI Generation to describe those in their formative educational years during the rise of Gen AI, students whose academic development is unfolding alongside the rapid adoption of these tools.
@@
 Their adaptability is being tested in real time.
-Students want more than access, they are asking for structure.
+Students want more than access; they are asking for structure.
 They seek guidance on using Gen AI responsibly, developing critical thinking around AI-generated outputs, and building a strong foundation in ethical reasoning.
@@
 Their voices point to an urgent need for Gen AI literacy that spans disciplines and prepares both students and faculty to engage with these tools thoughtfully and equitably.
-They envision a future where GenAI is not a source of confusion or inequity, but a tool for growth—integrated into education in ways that are ethical, empowering, and human-centered.
-6 List of abbreviations
-Abbreviation Full Form
-GenAI Generative Artificial Intelligence
-AI Artificial Intelligence
-IRB Institutional Review Board
-OMSCS Online Master of Science in Computer Science
-MSCS Master of Science in Computer Science
-OAE [Institution Name Redacted for Review] Office of Academic Effectiveness
-STEM Science, Technology, Engineering, and Mathematics
-CS Computer Science (used in tables/figures for grouping)
-7 Future Considerations
+They envision a future where Gen AI is not a source of confusion or inequity, but a tool for growth, integrated into education in ways that are ethical, empowering, and human-centered.
+VI Future Considerations
 To better understand the evolving impact of Gen AI on student development and educational outcomes, future research should consider longitudinal studies that track student perceptions, readiness, and concerns over time.
@@
 Capturing these changes across semesters or academic years would provide valuable insight into how exposure, policy, and pedagogy shape preparedness and perception in the long term.
-In addition to tracking perception, researchers should examine the cognitive consequences of increased reliance on GenAI tools, such as ChatGPT, for learning and problem-solving.
+In addition to tracking perception, researchers should examine the cognitive consequences of increased reliance on Gen AI tools for learning and problem-solving.
 Cognitive offloading may free students to focus on higher-order thinking, but it also raises questions about memory retention, deep learning, and critical reasoning.
-Understanding how sustained use of GenAI affects cognitive engagement, information retention, and academic identity will be critical to developing pedagogical strategies that support—not replace—core learning processes.
-There is also a need for targeted dissemination and outreach to disciplines beyond computer science.
-This study found that students in non-computer science fields, particularly those in the ”Other” category, demonstrated strong emotional responses and a desire for inclusion in GenAI discourse, despite limited curricular exposure.
-Expanding GenAI literacy and integration into these areas can help bridge the readiness gap and empower students across all domains to engage critically and confidently with this transformative technology.
-8 Declarations
-8.1 Funding
-This study was conducted without external funding. No funding agency was involved in the design, data collection, analysis, or publication of this research.
-8.2 Funding Statement
-8.3 Consent to Publish declaration
-Not applicable
-8.4 Consent to Participate
-This research study was Institutional Review Board (IRB) approved. Respondents provided informed consent.
-8.5 Ethics declaration
-This research study was Institutional Review Board (IRB) approved.
-The study collected students’ perceptions of their future career preparedness in light of the potential impact of generative artificial intelligence (GenAI) through a survey.
-The survey only collected responses for students located within the United States, and 18 years of age or older as per IRB instruction.
-Respondents provided informed consent.
-Survey questions were designed to protect participant identity.
-Participation was completely voluntary, and had the ability to opt out at any time.
-No benefit or compensation was provided.
-We complied with any applicable laws and regulations regarding confidentiality.
-The IRB and Office of Human Research Protections reserved the right to look at study records at any time.
-Contacts to the author and Office of Research Integrity Assurance were provided.
+Understanding how sustained use of Gen AI affects cognitive engagement, information retention, and academic identity will be critical to developing pedagogical strategies that support, not replace, core learning processes.
+Expanding Gen AI literacy and integration across disciplines can help bridge the readiness gap and empower students across all domains to engage critically and confidently with this transformative technology.
+Supplemental Materials
+Supplemental materials can be found under datasets at https://huggingface.co/sirajmicaela
+Acknowledgment
+Thank you to the Georgia Tech Office of Academic Effectiveness for their support in providing the undergraduate cohort sample and sample methodology.
+And thank you to the participants who shared their voice.